### v1

# The role of synthetic data in Multilingual, Multi-cultural AI systems: Lessons from Indic Languages

Abstract
Developing AI systems that operate effectively across languages while remaining culturally grounded is a long-standing challenge, particularly in low-resource settings. Synthetic data provides a promising avenue, yet its effectiveness in multilingual and multicultural contexts remains underexplored. We investigate the creation and impact of synthetic, culturally contextualized datasets for Indian languages through a bottom-up generation strategy that prompts large open-source LLMs ($\geq 235$B parameters) to ground data generation in language-specific Wikipedia content. This approach complements the dominant top-down paradigm of translating synthetic datasets from high-resource languages such as English. We introduce Updesh111[Uncaptioned image]https://hf.co/datasets/microsoft/Updesh_beta, a high-quality large-scale synthetic instruction-following dataset comprising 9.5M data points across 13 Indian languages, encompassing diverse reasoning and generative tasks with an emphasis on enhancing long-context and multi-turn capabilities, in addition to improving alignment with Indian cultural contexts. A comprehensive evaluation incorporating both automated metrics and human annotation across 10k assessments indicates that the generated data is of high quality; however, human evaluation highlights specific areas for further improvement. Additionally, we perform downstream evaluations by fine-tuning models on our dataset and assessing their performance across 15 diverse multilingual datasets to evaluate the generalizability of our dataset. Our experiments demonstrate that models trained on Updesh consistently obtain significant performance improvements on generative tasks and remain competitive on multiple-choice evaluations in NLU tasks. Further, the relative improvements of models fine-tuned with Updesh are most pronounced in low and medium-resource languages, effectively narrowing the gap between these languages and high-resource languages. These findings provide empirical evidence that effective multilingual AI development requires multi-faceted data curation and generation strategies that include context-aware, culturally grounded methodologies.
1 Introduction
Building multilingual and multicultural AI is essential to ensure equitable access to technology in diverse linguistic and cultural communities. Without this inclusivity, AI systems risk reinforcing global power imbalances by privileging dominant languages and Western-centric worldviews. However, even frontier language models often underperform in languages other than English and in non-Western contexts. This is not only due to limited linguistic and cultural diversity in pre-training data, but also to English-centric practices in fine-tuning, evaluation, and other parts of the development pipeline. While pre-training data is often treated as abundant and readily available due to large-scale web crawling, fine-tuning and evaluation data must be deliberately constructed and curated to ensure quality and relevance. This often leads to skipping the creation of truly multilingual and culturally grounded data altogether, or relying on simplified approaches like translating existing English datasets, often overlooking linguistic nuance and cultural context.
As highlighted by Joshi et al. (2020), there is a significant data gap between high-resource and low-resource languages in the internet data, the major source of pre-training data. Their taxonomy classifies world languages into six resource classes, with Classes 5-6 encompassing over $2400$ languages (93.8793.87% of all world languages) serving $1.2$ billion speakers, yet having minimal to no digital resources, creating a massive representation gap in current AI systems. This gap is even more pronounced for fine-tuning and evaluation datasets (Hu et al., 2025). Synthetic data generation has become popular to augment training sets and create benchmarks, by generating new data or by transforming existing data into desirable formats, such as converting raw text into structured instruction-style prompts. Synthetic data, created mainly in English, has shown promise in improving downstream performance in tasks such as reasoning (Goldie et al., 2025; Harsha et al., 2025), coding (Wei et al., 2024; Shao et al., 2025) and retrieval (Bonifacio et al., 2022; Dai et al., 2023; Chitale et al., 2025). In this work, we study whether synthetic data has the potential to supplement scarce multilingual, multicultural data when carefully designed.
Techniques for generating synthetic data are typically grounded in English-centric practices, often relying on assumptions about model quality and reliability that do not hold for many other languages. Assessing the quality and utility of synthetic data is an ongoing area of research. Current methods for evaluating quality include human and automated evaluation to assess fluency, correctness, and diversity, which may not suffice for assessing the quality of multilingual, multicultural synthetic data. Downstream evaluation involves fine-tuning models using synthetic data and evaluating models on standardized benchmarks. In this paper, we introduce the first framework for multilingual and multicultural synthetic data generation, focusing on key dimensions such as choice of generation strategies, grounding in language-specific data, quality assessment, and evaluation. We apply this framework to a case study that involves generating synthetic data across 13 Indian languages and conduct comprehensive evaluations to check for quality and utility. Our detailed analysis across multiple multilingual synthetic datasets offers practical insights for developing more inclusive, linguistically diverse, and culturally attuned language technologies.
2 Framework
Synthetic data is utilized throughout the AI lifecycle, including pre-training, supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF), and evaluation (Viswanathan et al., 2025). Synthetic data generation starts with a small set of human-curated seeds, which is then provided to an LLM to generate instructions, inputs and outputs, followed by filtering for correctness and quality. An alternate approach is to retrieve the relevant datasets and use an LLM to rewrite them into the desired format. The dominant method for producing multilingual synthetic data involves translating English data with Neural Machine Translation systems or Large Language Models (LLMs) for pre-training (Doshi et al., 2024; Khan et al., 2024), Supervised Fine Tuning (SFT) (Shaham et al., 2024), and evaluation. However, translation may introduce translation artifacts (referred to as translationese) (Zhang & Toral, 2019; Vanmassenhove et al., 2021) and yield weaker correlations with human evaluations compared to benchmarks tailored for particular languages and contexts (Kreutzer et al., 2025; Wu et al., 2025). Recent work has explored bottom-up approaches that rely on grounding and retrieval augmentation, including instruction backtranslation (Li et al., 2024) and instruction back-and-forth translation (Nguyen et al., 2024), which generate high-quality synthetic data grounded in web corpora. These methods can potentially demonstrate superior performance compared to traditional translation-based approaches. However, there remains limited exploration of such approaches in multilingual contexts. In this work, our focus is on curating multilingual, multicultural synthetic data for SFT, although our framework can apply to other kinds of multilingual, multicultural synthetic data, including synthetic pre-training data and preference data as well. Below, we describe the key factors to consider when generating this type of data.
Base model capability: Synthetic data is usually generated by prompting large foundation models due to their fluency, ability to follow instructions, and safety guardrails. For generating multilingual, multicultural synthetic data, candidate base models should be evaluated on established multilingual and multicultural benchmarks to determine performance in target languages. If benchmarks for a specific target language are unavailable, evaluation results from related languages or overall multilingual performance can be used as indicators. Some models list supported languages and may include multilingual benchmark results in their release reports, while others do not provide these details. Additional factors to consider when selecting a base model include licensing, cost, and whether models are open-weights or restricted.
Seed data selection: Seed data is typically selected on the basis of factors such as diversity and target task coverage. Additional considerations for multilingual, multicultural seed data collection would include prioritizing tasks with coverage of cultural knowledge, norms, and values, and tasks with more relevance in specific regional and cultural contexts (for example, agriculture).
Data generation strategy: Several data generation strategies can be considered for synthetic data generation, including translation, back-translation, and retrieval-augmented synthetic data generation. Translating high-quality, curated English SFT datasets allows us to carry over critical skills the LLM needs to acquire into the multilingual training data. Back-translation using existing multilingual datasets can enhance diversity. Retrieval-augmented generation from diverse sources can potentially leverage native speaker-authored content to produce synthetic data that includes cultural knowledge and linguistic nuances.
Quality metrics: Synthetic data should be evaluated for correctness and quality before being used in downstream tasks. This may involve both automated assessment and human evaluation by native speakers. Essential quality metrics for multilingual and multicultural synthetic data include:
  • Language correctness: Is the data in the right language, register, and dialect, or is there confusion with a similar language? (Marchisio et al., 2024)
  • Linguistic acceptability: Does the data sound fluent and natural to a native speaker of the language? (Hada et al., 2024b; a)
  • Cultural appropriateness: Does the data contain appropriate references to cultural artifacts and knowledge and accurately reflect values and norms?
  • Bias and safety: Does the data amplify stereotypes, contain unsafe content, or culturally inappropriate outputs?
Downstream evaluation The primary objective of generating synthetic data is to improve model performance in real-world applications, making the choice of downstream tasks for evaluations critical. This is particularly challenging given the scarcity of multilingual and multicultural benchmarks. To address this, downstream benchmarks should cover all target languages and, wherever possible, avoid relying on translated datasets originally in English. They should also include diverse tasks, especially when synthetic data spans multiple domains. They should incorporate tests of cultural knowledge, norms, and values. In some cases, new benchmarks aligned with the target task, language, and cultural context may need to be developed to accurately assess the utility of synthetic data. Finally, care must be taken to avoid benchmark contamination in the base model, a common issue with multilingual benchmarks (Ahuja et al., 2024).
Native speaker consent, involvement and privacy: Multilingual and multicultural synthetic data creation processes should involve native speakers in seed data selection and evaluation when feasible. Informed consent is required to address cultural considerations and data sovereignty according to local regulations. Synthetic data must not contain personally identifiable information derived from seed data or online sources.
3 Related Work
English Instruction Fine-tuning Datasets: Instruction Fine-tuning (IFT) (Ouyang et al., 2022) adapts pre-trained language models to follow human instructions through supervised learning on instruction-output pairs. English IFT datasets include the FLAN collection(Wei et al., 2022), which scaled instruction tuning to over 1,800 tasks and chain-of-thought prompting, and Self-Instruct (Wang et al., 2023), which demonstrated effectiveness of LLM-generated synthetic data, leading to Stanford Alpaca (Taori et al., 2023) and the enhanced Alpaca-GPT4 (Peng et al., 2023). Subsequent work expanded through conversational approaches like Vicuna (Chiang et al., 2023) using ShareGPT conversations, WizardLM’s Xu et al. (2024) Evol-Instruct methodology, and LIMA’s (Zhou et al., 2023a) demonstration that 1K curated prompts suffice for effective tuning. The ORCA series (Mukherjee et al., 2023; Mitra et al., 2023) introduced explanation tuning and Prompt Erasure techniques, culminating in OrcaAgent-Instruct (Mitra et al., 2024a) with 25.8M synthetic instruction-response pairs.
Multilingual Instruction Fine-tuning Datasets: Multilingual instruction-following capabilities have been developed through three primary methodologies. Translation-based methods create multilingual datasets by machine translation of existing English instruction data. Bactrian-X (Li et al., 2023) translates Alpaca (Taori et al., 2023) and Dolly (Conover et al., 2023) to produce 3.4M pairs across 52 languages, mAlpaca (Chen et al., 2024) demonstrates cost-effective multilingual tuning through Alpaca translation, and LIMA-X achieves 9.9% cross-lingual improvements using $\sim$1K instructions. Template-based approaches like M2Lingual leverage Evol-guided taxonomy (Xu et al., 2024) for synthetic generation across 70 languages. Hybrid strategies combine multiple techniques: Aya-Collection (Singh et al., 2024b) integrates crowdsourcing across 65 languages with templating and translation, repurposing xP3 (Muennighoff et al., 2023), Flan (Longpre et al., 2023), and Dolly datasets while employing NLLB 3.3B (Team et al., 2022) for n-way translation. Similarly, IndicAlign aggregates 74.7M prompt-response pairs across 20 Indian languages through aggregation of existing datasets, translations via the IndicTrans2 (Gala et al., 2023) model, synthetic conversation generation from India-centric Wikipedia, and crowdsourcing.
Data Generation Strategies: Most prior efforts rely on output distillation from advanced LLMs like GPT-4, where models learn from teacher-generated responses. However, recent approaches have explored alternative strategies to address distillation limitations. Instruction Backtranslation (Li et al., 2024) starts with a model fine-tuned on seed data to generate instruction prompts for web documents, followed by self-curation to select high-quality examples. This leverages web content diversity while avoiding over-reliance on teacher models through iterative self-augmentation and self-curation. Instruction Back-and-Forth Translation (Nguyen et al., 2024) extends backtranslation by generating synthetic instructions and then rewriting responses using an LLM to improve quality based on initial documents. This combines web information diversity with high-quality response generation, outperforming distillation and yielding better win rates than ShareGPT and Alpaca-GPT4 (Peng et al., 2023).
Limitations of prior multilingual IFT Datasets: Existing multilingual instruction datasets suffer from critical limitations that hinder effective cross-lingual alignment. Translation-based approaches like Bactrian-X, mAlpaca, and LIMA-X focus primarily on basic instruction-following while excluding advanced reasoning capabilities and lack demographic or cultural representation for local alignment. These datasets suffer from translation inaccuracies due to context-length limitations of traditional sentence-level translation methods (NLLB 3.3B and IndicTrans2) which fail to capture linguistic nuances and introduce errors that can potentially creep in during training. Despite impressive language coverage, Aya-Collection provides limited culturally-specific content, while IndicAlign sources a large proportion of data from WordNet (Miller, 1994) with a focus on question-answering offering limited task diversity. Most existing datasets are short-context and single-turn, failing to address real-world usage patterns including long-context or multi-turn dialogue capabilities. This can significantly constrain model performance in practical applications that demand sustained, contextually-aware interactions.
4 Data generation
The Updesh dataset comprises two complementary subsets addressing distinct aspects of multilingual instruction following: reasoning data and open-domain generative data. This design recognizes that reasoning capabilities are largely language- and culture-agnostic, making translation-based approaches suitable for tasks like mathematical problem-solving and logical inference (Shaham et al., 2024). Existing high-quality reasoning datasets such as OrcaAgent-Instruct and OrcaMath, thus are valuable resources for multilingual adaptation. However, generative capabilities requiring cultural awareness, linguistic naturalness, and factual grounding in local contexts cannot be adequately addressed through translation due to inherent Western-centric biases and lack of cultural specificity in existing datasets (Yao et al., 2024). Therefore, our generative subset employs a grounded approach that ensures factuality through Wikipedia-based content, maintaining linguistic naturalness through native language generation, and preserves cultural adherence through systematic curation of India-specific cultural artifacts.
Reasoning Data: Inspired by prior work (Ahuja et al., 2025; Khan et al., 2024), we translate eight subsets of the OrcaAgent-Instruct (Mitra et al., 2024a) and OrcaMath (Mitra et al., 2024b) datasets into 13 Indic languages. Specifically, we consider seven reasoning-related subsets from OrcaAgent-Instruct [^2] https://huggingface.co/datasets/microsoft/orca-agentinstruct-1M-v1 along with the Math subset from OrcaMath [^3] https://huggingface.co/datasets/microsoft/orca-math-word-problems-200k (Table 1). Both datasets have been attributed to induce significant chain-of-thought and reasoning capabilities in models during instruction-tuning without the need for specific preference optimization. We employ Llama-3.1-405B-Instruct for selective translation given its strong coverage in Indian languages and instruction-following capabilities that enable adaptation to various conversational styles (Sankar et al., 2025). Decoding is performed using nucleus sampling with top_p=0.95top\_p=0.95 and temperature set to $1.0$. After generation, all outputs undergo a series of quality control filters, described in detail below.
Table 1: Reasoning task categories and their descriptions
Task Type Description
Analytical Reasoning MCQ-style questions requiring step-by-step logical inference
Multiple-Choice Questions General-purpose problems across diverse knowledge domains
Fermi (Guesstimation) Open-ended estimation problems using logical assumptions
Few-Shot Chain-of-Thought Tasks with 4-5 in-context examples for learning
Brain Teasers Puzzles stimulating lateral thinking and creativity
Text Classification Categorization tasks for predefined labels
Math Grade-school arithmetic, algebra, and geometry word problems
Open-Domain Generative Data: Synthesizing generative data presents greater challenges than translating existing content, as it introduces increased risks of hallucinations, factual inaccuracies, and demographic misalignment. We conducted a comparative analysis between Llama-3.1-405B-Instruct and Qwen3-235B-A22B models across both reasoning and non-reasoning paradigms to evaluate qualitative differences in generated outputs. Manual assessment revealed that the Qwen3-235B-A22B model demonstrated superior proficiency in generative tasks and excelled in complex instruction following owing to its strong reasoning traces. Therefore, we used the Qwen3-235B-A22B for our generative data curation process. Drawing inspiration from instruction backtranslation techniques (Li et al., 2024), our methodology begins with unlabelled content to construct questions, followed by LLM-generated answers. To ensure generation diversity and contextual grounding that maintains factual accuracy and demographic relevance, we leverage Wikipedia pages in respective target languages as our knowledge base. Table 2 summarizes the eight generative task categories. As demonstrated, some tasks required two phases, which indicates two LLM inferences in order to generate the synthetic data instance.
Table 2: Generative task categories with synthesis methods, phases, and model configuration
Task Type Synthesis Method Phases Qwen3-Mode
Logical Reasoning Generate implicit inferences from text passages (1) Direct inference generation Reasoning
Multi-Hop QA Create questions requiring information synthesis across text segments (1) Question generation (2) Answer generation Reasoning
Creative Writing Transform factual content into engaging narratives (1) Generate creative piece (2) Generate eliciting prompt Reasoning
Multi-Turn Dialogue Agentic workflows with 3-5 turn conversations between personas (1) Generate dialog adhering to personas (2) Generate natural prompt Non-reasoning
Summarization Generate summaries preserving key information across languages (1) Direct summary generation Non-reasoning
Machine Translation Cross-lingual conversion maintaining cultural context (1) Direct translation Non-reasoning
Causal Reasoning Identify and explain cause-effect relationships in text (1) Direct causal analysis Reasoning
Further, to ensure cultural representation in synthetic data generation, we systematically curated culturally relevant content from Wikipedia using the MediaWiki API. Following (Yao et al., 2024)’s cultural taxonomy framework, we traversed Wikipedia category structures starting from Category:Culture of India and Category:Culture of India by state or union territory, exploring 2-3 levels deep. This yielded two complementary datasets: a general collection (54 categories, 19,143 artifacts) covering pan-Indian elements, and a state-wise collection (28 states + 8 union territories, 37,227 artifacts) with region-specific content spanning festivals, cuisine, traditional arts, architecture, and religious practices. We sampled 26.8K cultural artifacts from this collection to create multi-hop question-answer pairs for synthetic data generation.
Data Filtering After generating approximately 7.5 million data points across 13 languages for both the Reasoning and Open-Domain subsets, manual validation was not feasible, therefore, following the approach of Shen et al. (2025), we employed automated quality checks but use the standard threshold-based method instead of their anomaly detection-based method. Specifically, we applied two filtering criteria: (1) Language Identification using IndicLID (Madhani et al., 2023) with a 0.75 confidence threshold, and (2) word repetition ratio capped at 0.75 to flag low-quality generations. Table 3 shows the filtering results, demonstrating high data quality with drop rates below 2% for most subsets. The main exception is the FS-CoT subset for Urdu, where the outputs showed excessive repetition leading to higher filtering rates, but we maintain these thresholds to ensure data integrity. For the English-to-XX translation tasks, Assamese had the highest drop rate as the model frequently generated Bengali text instead, likely due to the shared script and similarity between these languages, and because Assamese is a low-resource language.
Table 3: Document filtering statistics averaged over all Indic Languages
Reasoning Subset Generative Subset
Category Total Drop Rate (%) Category Total Drop Rate (%)
analytical_reasoning 25,000 0.047 logical_reasoning 16,384 1.459
brain_teaser 50,000 0.043 multihop_QA 16,384 1.459
fermi 25,000 0.015 creative_writing 16,384 1.459
fs_cot_flow 25,000 3.769 Multi_turn_Dialog 16,384 1.611
math 200,035 0.035 summarization 16,384 1.526
mcq 99,986 0.135 machine_translation_en 16,384 0.641
rc 50,000 0.379 machine_translation_xx 16,384 17.047
text_classification 50,000 1.878 cultural_multihop_reasoning 26,834 0.347
causal_reasoning 16,384 1.453
5 Data Quality Analysis
5.1 Translation Quality Assessment for the Reasoning Subset
For the reasoning subset, we conducted large-scale selective translation utilizing the Llama-3.1-405B-Instruct. Given that the input often consists of long-context and non-standard text, it is crucial to rigorously evaluate the quality of the generated translations. To this end, following established methodologies the generated output in the target language is translated back into English and compared with the original English source. Due to the substantial volume of data, we randomly sample 4,096 instances from each subset for every language to ensure a representative evaluation. For the backtranslation step, we employed the more efficient Llama-3.3-70B-Instruct, which offers faster inference and serves as a conservative lower bound on translation quality. We assessed translation fidelity using standard metrics such as ChrF (Popović, 2015), implemented using the SACREBLEU library (Post, 2018) for consistency. Table 9 in the Appendix demonstrates consistently high backtranslation scores across all languages and subsets, indicating robust translation quality.
5.2 Data Quality Assessment for Generative Data
Although large language models (LLMs) have emerged as scalable evaluators following the popular LLM-as-a-judge paradigm, their reliability in culturally nuanced and low-resource language settings still remains limited  (Watts et al., 2024). Our study combined LLM-based assessment with native-speaker human annotation to quantify data quality and establish inter-annotator agreement. We employed a stratified sampling strategy to select 100 data points per task category: Creative Writing, Cultural Multi-Hop Reasoning, Multi-Turn Dialogue, and aggregated reasoning tasks (Logical, Causal, and Multi-Hop Reasoning) across five Indian languages (Assamese, Gujarati, Hindi, Malayalam and Punjabi) resulting in a total of 2000 data points. Our sampling methodology ensures representative coverage by partitioning responses into quintiles based on the output length and maintaining proportional sampling across buckets to preserve the original length distribution while capturing variance across responses. Human evaluations were conducted by native speakers through an external agency (Table 7). We employed GPT-4o as an automated evaluator using identical evaluation protocols to allow direct comparison. We devised the following metrics to capture multidimensional quality aspects:
  • Creative Writing: Instruction adherence, linguistic fluency, narrative coherence
  • Reasoning Tasks: Answer adequacy, contextual adherence, instruction compliance, fluency and readability, problematic content detection, cultural relevance assessment (cultural multi-hop only)
  • Multi-Turn Dialogue: Persona consistency, topical coherence, linguistic plausibility, repetition detection, toxicity screening, instruction adherence
  • Summarization: Content coverage, factual accuracy, conciseness, coherence and logical flow, stylistic appropriateness
  • Translation: Semantic correctness, fluency corectness, domain appropriateness, style and tone, completeness
Further details on the metrics and their definitions could be found in the Appendix in Table LABEL:tab:rubrics. All metrics utilize a three-point Likert scale (0-2) with detailed rubrics provided to both human and automated evaluators to ensure assessment consistency. The prompts used for the LLM evaluations could be found in the Appendix A.7. The guidelines given to the human annotators have been added in the supplementary material. Quantitative analysis of human evaluations reveals very high data quality: across 10,000 individual metric assessments, human evaluators assigned only 27 scores of zero (0.27%), indicating that the overwhelming majority of generated content meets acceptable quality thresholds. We plan to publicly release all evaluation data, including both human and GPT assessments, to ensure complete transparency and support future research on calibration of LLM-based evaluators.
Figure 1: Human LLM-judge agreement across evaluation metrics, revealing differences across dimensions.
Inter-Annotator Agreement: To quantify the reliability of automated evaluation, we computed the percentage agreement between the LLM-judge and the human evaluators. We find significant variance in agreement across metrics (Figure 1). Agreement substantially deteriorates for culturally and linguistically nuanced assessments, particularly for linguistic plausibility evaluation and in repetition detection in long dialogue sequences. Conversely, well-defined objective criteria, such as toxicity detection and problematic content identification, demonstrate robust inter-annotator alignment. Since our data is generated using benign prompts and topics, toxic or problematic content is expected to be very rare or nonexistent. These results are consistent with earlier research on the challenges of subjective versus objective evaluation and reveal shortcomings in current LLM-judges for assessing culturally sensitive content  (Watts et al., 2024). The distribution of scores across tasks and languages for both humans and LLM can be found in Appendix  A.2.
6 Downstream Tasks Evaluation
We selected two base models, Llama-3.1-8B and Phi4-14B, for fine-tuning experiments. These models were chosen based on their size (for feasibility of experiments given available resources) and reported multilingual capabilities (Grattafiori et al., 2024; Abdin et al., 2024). We used the Axolotl framework [^4] https://github.com/axolotl-ai-cloud/axolotl for all fine-tuning runs. Details regarding the hyperparameters and compute resources used can be found in Appendix A.4.
Baselines: We fine-tuned both Llama-3.1-8B and Phi4-14B on three high-quality, open-source IFT datasets that collectively cover nearly all the Indic languages included in our work: the Aya-Collection (Singh et al., 2024b), IndicAlign (Khan et al., 2024) and the Bactrian-X dataset (Li et al., 2023). To the best of our knowledge, these are the only open datasets that offer both broad language coverage and instruction-following format, and there is a lack of other large-scale open IFT datasets for Indian languages. The Bactrian-X dataset provides training samples for 10 out of our 13 target languages, excluding Assamese, Kannada and Punjabi. The Aya-Collection dataset includes all target languages except Punjabi. Since Aya-Collection contains millions of samples per language, we uniformly sub-sampled it to construct a balanced dataset of approximately 7 million samples which is comparable in scale to Updesh. For IndicAlign, the WordNet subset ($\sim$97M pairs) is disproportionately large, less diverse, and redundant, so we downsampled it to one instance per entry, yielding about 7.3M training pairs while retaining all other subsets.
Downstream Tasks: Our evaluation framework consists of three task categories to comprehensively assess model capabilities. Natural language understanding (NLU) tasks use multiple-choice questions to measure comprehension and reasoning through likelihood-based scoring. Natural language generation (NLG) tasks like translation and summarization, assess models’ ability to generate coherent and contextually appropriate outputs. For instruction-following evaluation, we translate established benchmarks (IFEval, IFBench) using GPT-4o (prompts in A.6), enabling assessment of both cross-lingual instruction-following (English responses to non-English instructions) and native language instruction-following. This design systematically identifies model strengths and weaknesses across diverse tasks, providing a holistic performance assessment. Dataset details are in Appendix A.5.
6.1 Results
Table 4: NLU task performance. All entries are accuracy (higher is better).
Model Setting MMLU MMLU-I MILU ARC-I BoolQ-I TVQA-I Bele INCL GMMLU
Llama Base 0.23 0.23 0.26 0.25 0.62 0.27 0.24 0.25 0.23
Bactrian 0.63 0.37 0.42 0.26 0.65 0.66 0.41 0.39 0.42
Indic-Align 0.60 0.34 0.42 0.27 0.71 0.59 0.32 0.34 0.39
Aya 0.59 0.33 0.39 0.28 0.76 0.51 0.47 0.34 0.37
Updesh 0.51 0.34 0.37 0.30 0.78 0.46 0.51 0.31 0.37
Phi4 Base 0.78 0.43 0.52 0.30 0.80 0.68 0.55 0.45 0.55
Bactrian 0.78 0.40 0.48 0.29 0.74 0.67 0.48 0.42 0.48
Indic-Align 0.77 0.45 0.51 0.32 0.76 0.62 0.52 0.45 0.54
Aya 0.78 0.45 0.53 0.33 0.73 0.72 0.61 0.48 0.53
Updesh 0.78 0.49 0.54 0.32 0.81 0.60 0.74 0.47 0.57
Performance on NLU Tasks: Table 4 shows the performance of both base models fine-tuned on all three IFT datasets, along with the performance of the base models on NLU tasks. Phi4-Updesh stands out as the most consistently strong configuration, attaining the best overall scores on MMLU-I, MILU, BoolQ-I, BeleBele, INCL, and GlobalMMLU. For the rest of the tasks, different models excel - in case of the Llama base model, Bactrian-X leads knowledge-intensive tasks, attaining the highest on MMLU and TVQA-I, while Updesh excels in instruction-style reasoning benchmarks (ARC-I, BoolQ-I) and performs best on reading comprehension tasks such as BeleBele. Phi4’s MMLU performance appears saturated across all fine-tuning baselines, suggesting a ceiling effect from pre-training knowledge or lack of coverage within all the training mixtures. Overall, these results demonstrate that different IFT datasets offer distinct advantages across various NLU capabilities.
Table 5: NLG task performance. All entries are ChrF scores (higher is better).
Model Setting XSum Flores En-XX Flores XX-En IN22-Conv-Doc
Llama Base 0.16 1.00 36.91 0.45
Bactrian 0.21 23.69 46.31 14.09
Indic-Align 9.26 25.18 2.29 10.98
Aya 0.24 22.60 0.53 6.73
Updesh 17.39 32.63 46.84 21.25
Phi4 Base 11.32 23.03 51.39 22.80
Bactrian 0.31 25.93 47.75 15.45
Indic-Align 0.27 27.48 1.77 13.82
Aya 0.37 26.51 7.63 14.01
Updesh 10.21 34.58 53.29 18.74
Performance on NLG Tasks: Table 5 demonstrates that Updesh is clearly superior to the other datasets on NLG tasks. Llama-Updesh emerges as the best configuration, achieving the highest performance across all generative tasks with substantial margins over all competing datasets as well as the base model. In case of Phi4, Updesh delivers the strongest translation performance, building effectively on the already capable base model. However, summarization results are more nuanced: the base model leads on XSum with Updesh being competitive, while the other fine-tuned models perform poorly. On long-context conversational translation (IN22-Conv-Doc) outperforms all fine-tuned variants, though it slightly lags behind the Phi4 Base model. Overall, we observe that the long-context generative training data in Updesh provides significant benefits for generative tasks.
Table 6: Instruction-following capabilities. All entries are accuracy (higher is better).
Model Setting mIFEval mIFBench
Native in English Native in English
Prompt Inst. Prompt Inst. Prompt Inst. Prompt Inst.
Llama Base 0.24 0.36 0.18 0.31 0.19 0.21 0.16 0.17
Bactrian 0.17 0.27 0.14 0.25 0.17 0.18 0.11 0.13
Indic-Align 0.16 0.24 0.14 0.23 0.17 0.19 0.16 0.17
Aya 0.13 0.22 0.11 0.21 0.18 0.19 0.18 0.19
Updesh 0.20 0.30 0.18 0.28 0.16 0.19 0.14 0.15
Phi4 Base 0.22 0.32 0.19 0.30 0.14 0.16 0.13 0.14
Bactrian 0.22 0.32 0.19 0.30 0.20 0.21 0.15 0.16
Indic-Align 0.16 0.25 0.15 0.24 0.15 0.17 0.14 0.16
Aya 0.15 0.25 0.14 0.24 0.20 0.21 0.19 0.20
Updesh 0.22 0.32 0.19 0.29 0.19 0.22 0.16 0.18
Performance on Instruction Following Tasks: Table 6 presents results on the GPT-4o translated multilingual versions of IFEval and IFBench, which evaluate fine-grained constraint adherence through highly specific requirements like character positioning, word counts, and formatting specifications - constraints notably absent in any training datasets. Native refers to prompts as well as expected output in regional language, while in English uses regional language prompts but expects English responses. Prompt-level evaluation measures whether all instructions in a prompt are followed, while instruction-level evaluation measures individual instruction adherence. Updesh emphasizes general instruction following involving task completion and output formatting, making these benchmarks an effective OOD evaluation for precise constraint adherence. For Llama, fine-tuning generally degrades performance, but Updesh shows smaller drops (mIFEval: 3-4 point lesser degradation than next best - Bactrian; mIFBench: comparable or better performance than other baselines, lags Aya on in English tasks). Phi4 exhibits different dynamics - Updesh achieves positive gains on mIFBench over the Base model (+4/+2 points Native; +1/+3 points English) while maintaining base-level performance on mIFEval, outperforming other fine-tuning approaches (Aya, Indic-Align) that show mixed or negative changes. Updesh provides robust instruction-following with minimal catastrophic forgetting compared to other baselines.
Language-wise trends: Figure 2 shows a clear trend: performance improves as language resources increase for both models and task types (NLU accuracy and NLG ChrF). We illustrate this with the resource taxonomy Joshi et al. (2020), ordered from 0 (lowest-resource) to 5 (high resource). The relative improvements with Updesh are most pronounced in low/mid-resource bands, effectively narrowing the gap to high-resource languages. The base model (without fine-tuning) is the weakest, with Bactrian-X and IndicAlign often competitive in the medium resource languages while Aya-Collection is more variable. Notably, generation benefits more from data quality: NLG improvements increase more dramatically as languages become lower-resourced, most clearly for Phi4-Updesh. Across language-resource tiers, Updesh is consistently strongest, especially at higher resource levels, reaching about $\sim$0.65 for Llama (NLU), $\sim$0.76 for Phi-4 (NLU) in terms of average accuracy, and the highest ChrF for Phi-4 (NLG).
Figure 2: NLU and NLG task performance grouped by language–resource class taxonomy from Joshi et al. (2020). Updesh yields the largest relative gains in low/mid-resource languages
6.2 Discussion
Quality evaluation: Our comprehensive quality evaluation ranges from basic filters to nuanced human evaluation. Experiments with LLM-based evaluators show that while LLM-judges do well on some metrics, they struggle with subtle multilingual and multicultural contexts. For example, they often misjudge how a language is typically spoken or written, and tend to lost track of persona consistency in longer dialogues. These issues motivate the need for scalable and reliable evaluation methods that go beyond surface-level checks.
Task-type sensitivity (NLU vs.NLG): We observe variation by task type - for instance, on NLU-style tasks (usually multiple-choice), no configuration is a clear winner across all datasets; performance depends on the base model and the specific benchmark. In contrast, on NLG-style tasks that score free-form generations, models trained on Updesh consistently perform strongly. This pattern suggests that the quality control and discourse control learned during training is transferred more directly to NLG than to choice-based NLU.
Evaluation protocol differences matter: Most NLU evaluations compute correctness via likelihood comparisons over (context, option) pairs, effectively testing discrimination among short alternatives. NLG evaluations, by design, require longer generations and measure faithfulness, adequacy, and fluency under open-ended prompting. The divergence in scoring objectives partially explains why a dataset that emphasizes long-form instruction-following (like Updesh) yields larger gains on NLG than on likelihood-based NLU. We find that the Bactrian dataset tends to perform best on NLU-style benchmarks, which can be attributed to format alignment and a larger proportion of MCQ or short-from prompts.
Distribution mismatches between train and test: Updesh contains longer-context samples and lacks short-text MCQ-style items; its task composition emphasizes multi-step, generative instruction-following. Even the MCQ subset which is part of the reasoning data in Updesh consists of long-form answers to MCQs, with close to no representation of short-text or direct answers. In contrast, many NLU benchmarks, however, are MCQ-heavy and short-context. This composition shift (task format, length, and response style) likely reduces the relative NLU gains for Updesh-trained models while amplifying NLG gains, showing that models generalize best to downstream tasks with evaluation formats that resemble their training distribution.
7 Conclusion
In this work, we examined synthetic data as a potential remedy for the scarcity of multilingual and multicultural resources. Through a comprehensive framework and systematic experiments across Indian languages, we identified effective strategies for data generation, quality assessment, and downstream evaluation, beyond English-centric norms. We built Updesh, a 9.5M 13-language IFT dataset using a culturally grounded, bottom-up pipeline. Our comprehensive evaluation spanning data generation, quality assessment (human and LLM-as-judge), and downstream tasks, revealed that synthetic data can potentially bridge resource gaps. However, no single recipe dominates across tasks and models, motivating approaches tailored to diverse linguistic and cultural contexts. We will release the Updesh dataset, evaluation protocols, and detailed analyses to enable transparent, reproducible progress.
8 Ethical Considerations
Our discussion of ethical considerations is guided by the framework proposed by Bender & Friedman (2018).
Institutional Process and Oversight
The data annotation was conducted by a third-party vendor and was approved by the Institutional Review Board of our organization and by the vendor.
Data Provenance and Quality Assurance
To mitigate potential artifacts and quality issues in the synthetic data, we implemented a rigorous quality control process. This process involved both automated evaluation with GPT-4o and manual verification by human annotators. We observed a high concordance between automated and human judgments on potentially problematic content. Furthermore, given that the data was generated using state-of-the-art large language models, the baseline incidence of such content was already substantially reduced. In a manual evaluation of 500 samples, human annotators flagged only 1 sample (0.2%) on metrics pertaining to problematic content, confirming the high quality of the resulting dataset.
Annotator Demographics
Annotators were recruited through a professional external services company. All annotators assigned to a given data point were native speakers of the language represented in the data. Table 7 summarizes the annotator demographics (education, region, age distribution, and gender). Each data worker was compensated at the rate of \$2 per data point.
Category Summary
Participants 15
Qualification Post-graduation: 7
Graduation: 8
Geography Spread across 8 Indian states
Age distribution 21–30: 7
31–40: 5
41–50: 3
Gender Female: 11
Male: 4
Table 7: Participant demographics summary
Appendix A Appendix
A.1 Rubrics used for the quality evaluation of the synthetic data
Table 8: Evaluation Rubrics for AI Tasks
Task Rubric Description
Creative Writing Instruction adherence Assesses if the output strictly follows all constraints and guidelines provided in the prompt.
Fluency Evaluates the naturalness, grammatical correctness, and readability of the generated text.
Narrative coherence Checks for logical consistency in the plot, character development, and thematic elements.
Reasoning Tasks Answer adequacy Determines if the final answer is correct, complete, and directly addresses the core question.
Context adherence Measures whether the reasoning remains faithful to the provided context, avoiding external facts.
Instruction adherence Verifies that the output’s structure, format, and steps match the user’s instructions.
Fluency and readability Assesses the clarity, logical flow, and ease of understanding of the explanation.
Problematic content and cultural relevance Scrutinizes the response for harmful stereotypes and ensures it is culturally sensitive and appropriate.
Multi-turn Dialog Persona Adherence Evaluates the model’s ability to consistently maintain a specific character or role throughout the conversation.
Topic adherence Checks if the conversation remains focused on the established topic or transitions logically.
Linguistic plausibility Assesses whether the dialogue sounds natural, human-like, and contextually appropriate.
Repetitiveness Measures the degree to which the model avoids unnecessarily repeating phrases or ideas.
Toxicity check Ensures the response is free from any offensive, harmful, or inappropriate content.
Instruction adherence Verifies that the model follows meta-instructions given by the user during the dialogue.
Summarization Coverage Determines if the summary successfully captures all essential points from the source text.
Factual accuracy Checks that the summary correctly represents the information and facts from the original document.
Conciseness Evaluates whether the summary is significantly shorter than the source while retaining critical information.
Coherence and logical flow Assesses if the summary is well-structured, logically organized, and easy to follow.
Style and tone Measures how well the summary reflects the style and tone of the original text.
Translation Semantic correctness Assesses whether the meaning, intent, and nuance of the source text are accurately conveyed.
Fluency correctness Evaluates the grammatical accuracy and naturalness of the translated text in the target language.
Domain appropriateness Checks if the terminology is correct and suitable for the specific subject matter (e.g., legal, medical).
Style and tone Determines if the translation successfully captures the original author’s writing style and emotional tone.
Completeness Verifies that the entire source text has been translated without any omissions or additions.
A.2 Results from the quality evaluations of the synthetic data
Figure 3: LLM evaluations across 5 synthetically generated tasks
Figure 4: Expert human evaluations across 5 synthetically generated tasks
Figures  3 and 4 show the distribution of the scores received for the tasks we evaluated across the languages for the tasks. Expert human evaluators have consistently given a score of 2 across languages and tasks, indicating the high quality of Updesh. The disagreements between humans and the LLM is however unclear from these plots. We thereby, performed a thorough inter annotator analysis, the details of which are present in the next figures.
Figure 5: Confusion matrices showing agreement between human and LLM evaluators
Figure 6: Agreement between human and LLM evaluators per task and language respectively
Continuing the claims made from Figure 1, Figure 5 provides a clearer view of the tasks on which humans and LLMs are likely to agree or disagree. We find that the largest disagreements occur in tasks such as assessing the linguistic plausibility of a given text in a regional language. Furthermore, LLMs struggle with evaluating long-context tasks, such as evaluating whether the same persona is maintained throughout a multi-turn conversation in a regional language. There is also a notable divergence between what human evaluators consider fluent in a relatively low-resource language and what an LLM deems fluent. In contrast, we observe considerable agreement in tasks like toxicity detection and problematic content flagging. LLMs also perform reasonably well at identifying whether a text is culturally relevant, but issues arise when the evaluation requires more fine-grained judgments of multilinguality and multiculturalism. We do not identify any specific trends language or task-wise as apparent in Figure6.
A.3 Data Quality Assesment of Reasoning Data
Language analytical brain fermi fs (cot_flow) math mcq rc text_class
Assamese 75.02 71.62 79.93 96.59 79.87 64.33 65.07 71.29
Bengali 87.25 77.40 80.69 82.10 79.87 67.25 74.94 74.72
Gujarati 77.86 67.15 82.14 73.96 49.66 78.04 63.21 55.95
Hindi 84.49 79.23 81.54 87.11 64.25 74.80 73.71 67.00
Kannada 79.96 76.87 80.02 81.26 65.80 69.91 64.15 60.77
Malayalam 75.21 73.41 70.10 77.93 68.51 55.69 63.19 75.63
Marathi 77.84 68.63 69.33 82.81 68.56 64.48 56.49 60.58
Nepali 81.79 86.18 74.71 83.32 53.98 56.42 53.86 59.63
Odia 56.47 62.06 50.70 93.61 57.55 51.21 42.95 52.15
Punjabi 83.75 51.79 77.40 79.04 61.19 51.21 69.81 54.11
Tamil 79.28 70.70 74.54 75.16 57.83 60.66 63.94 49.65
Telugu 78.24 80.26 74.33 79.91 69.88 60.66 61.93 60.95
Urdu 85.05 79.97 66.31 81.93 59.80 61.76 64.91 67.23
Table 9: Backtranslation ChrF scores for the Reasoning subset.
A.4 Hyperparameters
Table 10: Training hyperparameters used for all our experiments
Hyperparameter Value
Base Model phi4-base-hf / llama-3.1-8b-hf
Sequence Length 65,536
Effective Batch Size 8192
Number of Epochs 3
Optimizer AdamW Torch Fused
Learning Rate 1.0e-05
LR Scheduler Cosine
Adam Beta1 0.9
Adam Beta2 0.95
Max Grad Norm 1.0
Warmup Ratio 0.03
Weight Decay 0.1
NEFTune Noise Alpha 5
Precision BF16
Flash Attention True
Gradient Checkpointing True
A.5 Evaluation Datasets
Table 11: Evaluation datasets
Dataset Source
NLU Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021)
MMLU Indic (MMLU-I) SarvamAI
ARC Indic (ARC-I) SarvamAI
BoolQ Indic (BoolQ-I) SarvamAI
TriviaQA Indic (TVQA-I) SarvamAI
BeleBele (Bele) Bandarkar et al. (2024)
INCLUDE (INCL) Romanou et al. (2025)
Global MMLU (GMMLU) Singh et al. (2025)
NLG Extreme Summarization (Xsum) Singh et al. (2024a)
Flores English to Other (Flores EnXX) Goyal et al. (2022)
Flores XX-En (Flores XXEn)
IN22-Conv (in22-conv-doc) Gala et al. (2023); Gumma et al. (2025)
IF Multilingual Instruction Following Evaluation (mIFEval) Zhou et al. (2023b)
Multilingual Instruction Following Benchmark (mIFBench) Pyatkin et al. (2025)
A.6 Prompt used for conversion of IFEval and IFBench
Translation Generation Prompt INSTRUCTION: You are a professional translator. Your sole task is to translate the provided English text into {target_language}. Rules: Output only the translation—nothing else (no explanations, no commentary). Do not answer or solve any task contained in the text; just translate it. Preserve formatting, punctuation, numbers, placeholders, and bracketed tokens exactly as they appear. Translate idiomatically and naturally for a fluent native speaker of {target_language}. If the input is an instruction (e.g., “Write a resume for …”), translate the instruction itself—do not perform it. Do not introduce extra formatting (e.g., don’t create JSON blocks unless the source already has them). SOURCE LANGUAGE: English TARGET LANGUAGE:
{target_language}
TEXT TO TRANSLATE:
{source_text}
OUTPUT REQUIREMENT:
Return only the translated text in {target_language} with formatting faithfully preserved.
A.7 Sample Prompts Used for Evaluation
The following are representative prompts (one per task) used to evaluate different task types. Each prompt follows the established rubrics from Table LABEL:tab:rubrics.
A.7.1 Creative Writing - Instruction Adherence
Creative Writing Instruction Adherence Evaluation Prompt INSTRUCTION: You are an expert literary critic and evaluator, tasked with assessing the degree to which a synthetically generated creative piece adheres to the user’s writing prompt. You are required to read the USER QUESTION thoroughly and analyze how well the generated response incorporates the specified narrative elements, stylistic choices, and constraints. USER QUESTION: {user_prompt} TARGET LANGUAGE: {tgt_lang} ASSISTANT GENERATED RESPONSE: {assistant_output} A response that adheres to the user’s creative brief has: Narrative & Thematic Completeness: It fully incorporates all requested characters, plot points, settings, and themes. The creative piece feels complete and resolves according to the prompt’s guidelines. Stylistic & Tonal Adherence: The content’s tone, mood, and writing style (e.g., genre conventions, a specific author’s voice) directly match the user’s request. Format & Constraint Compliance: It follows all explicit formatting requirements (e.g., poem, script, short story) and abides by any constraints (e.g., word count, inclusion/exclusion of specific words, use of certain literary devices). Creative Intent Alignment: It successfully captures the spirit and intended artistic goal of the user’s prompt, creating a piece that feels like a faithful realization of the user’s idea. Use the following scoring scale: 5 – Excellent: The response masterfully incorporates all creative constraints, including plot, character, tone, style, and format. It not only follows the instructions to the letter but also demonstrates a creative flair that enhances the user’s original idea. 4 – Good: The response successfully incorporates most creative instructions. There may be minor deviations in tone or style, or a secondary plot/character element might be slightly underdeveloped, but the core creative vision is clearly and effectively realized. 3 – Fair: The response addresses some of the key creative instructions but neglects or misinterprets others. For instance, it might follow the plot but fail to capture the requested tone, or it might ignore a crucial character trait or constraint. 2 – Poor: The response shows significant deviation from the creative brief. It may follow a single, simple instruction (like the general topic) but disregards crucial constraints like genre, character personality, plot structure, or mood. 1 – Unacceptable: The response completely disregards the creative instructions. The generated text is thematically, structurally, and stylistically unrelated to the user’s prompt. Return your evaluation in the following JSON format:
  "score": <integer from 1 to 5>,
  "reason": "<brief explanation for the score>"
Do not include markdown, comments, or anything outside the JSON.
A.7.2 Multi-turn Dialog - Persona Adherence
Multi-turn Dialog Evaluation Prompt INSTRUCTION: You are an expert evaluator tasked with assessing how well a multi-turn dialog maintains persona consistency. You must analyze the conversation to determine if the assistant consistently embodies the specified character or role throughout the interaction. USER QUESTION: {user_prompt} TARGET LANGUAGE: {tgt_lang} ASSISTANT GENERATED RESPONSE: {assistant_output} A response that demonstrates strong persona adherence has: Character Consistency: The assistant maintains the same personality traits, speaking style, and behavioral patterns throughout the conversation. Role-Appropriate Knowledge: The responses reflect knowledge and expertise appropriate to the specified persona. Consistent Voice: The tone, vocabulary, and manner of speaking remain true to the character across all turns. Believable Interactions: The persona feels authentic and natural in the conversational context. Use the following scoring scale: 5 – Excellent: Perfect persona consistency with natural, believable character embodiment throughout all turns. 4 – Good: Strong persona adherence with minor inconsistencies that don’t break character immersion. 3 – Fair: Generally maintains persona but has noticeable lapses or inconsistencies in character. 2 – Poor: Significant persona inconsistencies that frequently break character immersion. 1 – Unacceptable: Complete failure to maintain persona or embody the specified character. Return your evaluation in the following JSON format:
  "score": <integer from 1 to 5>,
  "reason": "<brief explanation for the score>"
Do not include markdown, comments, or anything outside the JSON.
A.7.3 Reasoning Tasks - Answer Adequacy
Reasoning Task Evaluation Prompt INSTRUCTION: You are an expert evaluator specializing in logical reasoning and problem-solving. Your task is to assess whether the generated response provides a correct, complete, and well-reasoned answer to the given question. USER QUESTION: {user_prompt} TARGET LANGUAGE: {tgt_lang} ASSISTANT GENERATED RESPONSE: {assistant_output} A response with excellent answer adequacy demonstrates: Correctness: The final answer is factually accurate and logically sound. Completeness: All aspects of the question are addressed without omitting important elements. Direct Relevance: The response directly answers what was asked without unnecessary tangents. Clear Reasoning: The logical steps leading to the conclusion are evident and valid. Use the following scoring scale: 5 – Excellent: Completely correct and comprehensive answer with clear, logical reasoning. 4 – Good: Correct answer with minor gaps in completeness or explanation clarity. 3 – Fair: Generally correct but missing some important aspects or contains minor errors. 2 – Poor: Partially correct but has significant errors or omissions in reasoning or conclusion. 1 – Unacceptable: Incorrect answer or completely fails to address the question asked. Return your evaluation in the following JSON format:
  "score": <integer from 1 to 5>,
  "reason": "<brief explanation for the score>"
Do not include markdown, comments, or anything outside the JSON.

### v2

@@
-The role of synthetic data in Multilingual, Multi-cultural AI systems: Lessons from Indic Languages
+UPDESH: Synthesizing Grounded Instruction Tuning Data for 13 Indic Languages
 Abstract
-Developing AI systems that operate effectively across languages while remaining culturally grounded is a long-standing challenge, particularly in low-resource settings.
-Synthetic data provides a promising avenue, yet its effectiveness in multilingual and multicultural contexts remains underexplored.
-We investigate the creation and impact of synthetic, culturally contextualized datasets for Indian languages through a bottom-up generation strategy that prompts large open-source LLMs ($\geq 235$B parameters) to ground data generation in language-specific Wikipedia content.
-This approach complements the dominant top-down paradigm of translating synthetic datasets from high-resource languages such as English.
-We introduce Updesh111[Uncaptioned image]https://hf.co/datasets/microsoft/Updesh_beta, a high-quality large-scale synthetic instruction-following dataset comprising 9.5M data points across 13 Indian languages, encompassing diverse reasoning and generative tasks with an emphasis on enhancing long-context and multi-turn capabilities, in addition to improving alignment with Indian cultural contexts.
-A comprehensive evaluation incorporating both automated metrics and human annotation across 10k assessments indicates that the generated data is of high quality; however, human evaluation highlights specific areas for further improvement.
-Additionally, we perform downstream evaluations by fine-tuning models on our dataset and assessing their performance across 15 diverse multilingual datasets to evaluate the generalizability of our dataset.
-Our experiments demonstrate that models trained on Updesh consistently obtain significant performance improvements on generative tasks and remain competitive on multiple-choice evaluations in NLU tasks.
-Further, the relative improvements of models fine-tuned with Updesh are most pronounced in low and medium-resource languages, effectively narrowing the gap between these languages and high-resource languages.
-These findings provide empirical evidence that effective multilingual AI development requires multi-faceted data curation and generation strategies that include context-aware, culturally grounded methodologies.
+Developing culturally grounded multilingual AI systems remains challenging, particularly for low-resource languages.
+While synthetic data offers promise, its effectiveness in multilingual and multicultural contexts is underexplored.
+We investigate bottom-up synthetic data generation using large open-source LLMs ($\geq 235$B parameters) grounded in language-specific Wikipedia content, complementing dominant top-down translation-based approaches from English.
+We introduce Updesh, a high-quality large-scale synthetic instruction-following dataset comprising 9.5M data points across 13 Indian languages and English, encompassing diverse reasoning and generative tasks.
+Comprehensive evaluation using automated metrics and 10K human assessments confirms high data quality.
+Downstream evaluations performed by fine-tuning models on various datasets and assessing performance across 13 diverse multilingual datasets and model comparative evaluations, demonstrate that models trained on Updesh consistently obtain significant improvements on NLU, NLG evaluations.
+Finally, through ablation studies and cultural evaluations, we show that context-aware, culturally grounded data generation is essential for effective multilingual AI development .
+UPDESH: Synthesizing Grounded Instruction Tuning Data for 13 Indic Languages
+Pranjal A.
+Chitale   Varun Gummathanks: Work done at Microsoft   Sanchit Ahuja11footnotemark: 1   Prashant Kodali Manan Uppadhyay   Deepthi Sudharsan11footnotemark: 1   Sunayana Sitaramthanks: Corresponding author Microsoft Corporation   Nanyang Technological University   Northeastern University   Independent Researcher pranjalchitale@gmail.com, sunayana.sitaram@microsoft.com
 1 Introduction
-Building multilingual and multicultural AI is essential to ensure equitable access to technology in diverse linguistic and cultural communities.
-Without this inclusivity, AI systems risk reinforcing global power imbalances by privileging dominant languages and Western-centric worldviews.
-However, even frontier language models often underperform in languages other than English and in non-Western contexts.
-This is not only due to limited linguistic and cultural diversity in pre-training data, but also to English-centric practices in fine-tuning, evaluation, and other parts of the development pipeline.
-While pre-training data is often treated as abundant and readily available due to large-scale web crawling, fine-tuning and evaluation data must be deliberately constructed and curated to ensure quality and relevance.
-This often leads to skipping the creation of truly multilingual and culturally grounded data altogether, or relying on simplified approaches like translating existing English datasets, often overlooking linguistic nuance and cultural context.
-As highlighted by Joshi et al.
-(2020), there is a significant data gap between high-resource and low-resource languages in the internet data, the major source of pre-training data.
-Their taxonomy classifies world languages into six resource classes, with Classes 5-6 encompassing over $2400$ languages (93.8793.87% of all world languages) serving $1.2$ billion speakers, yet having minimal to no digital resources, creating a massive representation gap in current AI systems.
-This gap is even more pronounced for fine-tuning and evaluation datasets (Hu et al., 2025).
-Synthetic data generation has become popular to augment training sets and create benchmarks, by generating new data or by transforming existing data into desirable formats, such as converting raw text into structured instruction-style prompts.
-Synthetic data, created mainly in English, has shown promise in improving downstream performance in tasks such as reasoning (Goldie et al., 2025; Harsha et al., 2025), coding (Wei et al., 2024; Shao et al., 2025) and retrieval (Bonifacio et al., 2022; Dai et al., 2023; Chitale et al., 2025).
-In this work, we study whether synthetic data has the potential to supplement scarce multilingual, multicultural data when carefully designed.
-Techniques for generating synthetic data are typically grounded in English-centric practices, often relying on assumptions about model quality and reliability that do not hold for many other languages.
-Assessing the quality and utility of synthetic data is an ongoing area of research.
-Current methods for evaluating quality include human and automated evaluation to assess fluency, correctness, and diversity, which may not suffice for assessing the quality of multilingual, multicultural synthetic data.
-Downstream evaluation involves fine-tuning models using synthetic data and evaluating models on standardized benchmarks.
-In this paper, we introduce the first framework for multilingual and multicultural synthetic data generation, focusing on key dimensions such as choice of generation strategies, grounding in language-specific data, quality assessment, and evaluation.
-We apply this framework to a case study that involves generating synthetic data across 13 Indian languages and conduct comprehensive evaluations to check for quality and utility.
-Our detailed analysis across multiple multilingual synthetic datasets offers practical insights for developing more inclusive, linguistically diverse, and culturally attuned language technologies.
-2 Framework
-Synthetic data is utilized throughout the AI lifecycle, including pre-training, supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF), and evaluation (Viswanathan et al., 2025).
-Synthetic data generation starts with a small set of human-curated seeds, which is then provided to an LLM to generate instructions, inputs and outputs, followed by filtering for correctness and quality.
-An alternate approach is to retrieve the relevant datasets and use an LLM to rewrite them into the desired format.
-The dominant method for producing multilingual synthetic data involves translating English data with Neural Machine Translation systems or Large Language Models (LLMs) for pre-training (Doshi et al., 2024; Khan et al., 2024), Supervised Fine Tuning (SFT) (Shaham et al., 2024), and evaluation.
-However, translation may introduce translation artifacts (referred to as translationese) (Zhang & Toral, 2019; Vanmassenhove et al., 2021) and yield weaker correlations with human evaluations compared to benchmarks tailored for particular languages and contexts (Kreutzer et al., 2025; Wu et al., 2025).
-Recent work has explored bottom-up approaches that rely on grounding and retrieval augmentation, including instruction backtranslation (Li et al., 2024) and instruction back-and-forth translation (Nguyen et al., 2024), which generate high-quality synthetic data grounded in web corpora.
-These methods can potentially demonstrate superior performance compared to traditional translation-based approaches.
-However, there remains limited exploration of such approaches in multilingual contexts.
-In this work, our focus is on curating multilingual, multicultural synthetic data for SFT, although our framework can apply to other kinds of multilingual, multicultural synthetic data, including synthetic pre-training data and preference data as well.
-Below, we describe the key factors to consider when generating this type of data.
-Base model capability: Synthetic data is usually generated by prompting large foundation models due to their fluency, ability to follow instructions, and safety guardrails.
-For generating multilingual, multicultural synthetic data, candidate base models should be evaluated on established multilingual and multicultural benchmarks to determine performance in target languages.
-If benchmarks for a specific target language are unavailable, evaluation results from related languages or overall multilingual performance can be used as indicators.
-Some models list supported languages and may include multilingual benchmark results in their release reports, while others do not provide these details.
-Additional factors to consider when selecting a base model include licensing, cost, and whether models are open-weights or restricted.
-Seed data selection: Seed data is typically selected on the basis of factors such as diversity and target task coverage.
-Additional considerations for multilingual, multicultural seed data collection would include prioritizing tasks with coverage of cultural knowledge, norms, and values, and tasks with more relevance in specific regional and cultural contexts (for example, agriculture).
-Data generation strategy: Several data generation strategies can be considered for synthetic data generation, including translation, back-translation, and retrieval-augmented synthetic data generation.
-Translating high-quality, curated English SFT datasets allows us to carry over critical skills the LLM needs to acquire into the multilingual training data.
-Back-translation using existing multilingual datasets can enhance diversity.
-Retrieval-augmented generation from diverse sources can potentially leverage native speaker-authored content to produce synthetic data that includes cultural knowledge and linguistic nuances.
-Quality metrics: Synthetic data should be evaluated for correctness and quality before being used in downstream tasks.
-This may involve both automated assessment and human evaluation by native speakers.
-Essential quality metrics for multilingual and multicultural synthetic data include:
-  • Language correctness: Is the data in the right language, register, and dialect, or is there confusion with a similar language? (Marchisio et al., 2024)
-  • Linguistic acceptability: Does the data sound fluent and natural to a native speaker of the language? (Hada et al., 2024b; a)
-  • Cultural appropriateness: Does the data contain appropriate references to cultural artifacts and knowledge and accurately reflect values and norms?
-  • Bias and safety: Does the data amplify stereotypes, contain unsafe content, or culturally inappropriate outputs?
-Downstream evaluation The primary objective of generating synthetic data is to improve model performance in real-world applications, making the choice of downstream tasks for evaluations critical.
-This is particularly challenging given the scarcity of multilingual and multicultural benchmarks.
-To address this, downstream benchmarks should cover all target languages and, wherever possible, avoid relying on translated datasets originally in English.
-They should also include diverse tasks, especially when synthetic data spans multiple domains.
-They should incorporate tests of cultural knowledge, norms, and values.
-In some cases, new benchmarks aligned with the target task, language, and cultural context may need to be developed to accurately assess the utility of synthetic data.
-Finally, care must be taken to avoid benchmark contamination in the base model, a common issue with multilingual benchmarks (Ahuja et al., 2024).
-Native speaker consent, involvement and privacy: Multilingual and multicultural synthetic data creation processes should involve native speakers in seed data selection and evaluation when feasible.
-Informed consent is required to address cultural considerations and data sovereignty according to local regulations.
-Synthetic data must not contain personally identifiable information derived from seed data or online sources.
+Building multilingual, multicultural AI is essential for equitable access across communities.
+Yet frontier models often underperform in non-English and non-Western settings because diversity is limited in pre-training corpora and English-centric choices pervade the development pipeline (Joshi et al., 2020).
+While large-scale crawling can expand pre-training data, fine-tuning and evaluation sets require deliberate construction; translation-only approaches often overlook linguistic nuance and cultural context.
+Joshi et al.
+(2020) identify a stark imbalance in web-scale pretraining data: their lowest-resource categories (Classes 5–6) cover over 2.4K languages (93.87% of the world’s languages) spoken by 1.2B people, yet remain severely underrepresented online.
+This gap is more pronounced for fine-tuning and evaluation datasets (Hu et al., 2025).
+Synthetic data shows promise in English for reasoning (Goldie et al., 2025; Harsha et al., 2025), coding (Wei et al., 2024; Shao et al., 2025), and retrieval (Bonifacio et al., 2022; Dai et al., 2023; Chitale et al., 2025), but pipelines often embed English-centric quality assumptions that may not transfer.
+Evaluating multilingual, multicultural synthetic data thus remains open: standard human and automatic checks (fluency, correctness, diversity) are often insufficient, and downstream utility must be validated via fine-tuning and benchmarking.
+Figure 1: Overview of the data generation pipeline for the Updesh Dataset.
+In this work, we introduce Updesh, a culturally grounded multilingual synthetic instruction dataset with 9.5M examples, and provide extensive analyses of its quality and downstream utility.
+Models fine-tuned on Updesh consistently improve NLG and NLU performance, across languages and shows a considerable uplift in cultural evaluations as well.
+Alongside the dataset, we distill a set of design considerations for multilingual and multicultural synthetic data generation spanning generation strategies, language-specific grounding, quality assessment, and evaluation.
+Our results also expose limitations of LLM-as-judge in culturally nuanced settings and yield actionable guidance for dataset design, filtering, and downstream validation, enabling reproducible alternatives to translation-first pipelines.
 2 Related Work
-English Instruction Fine-tuning Datasets: Instruction Fine-tuning (IFT) (Ouyang et al., 2022) adapts pre-trained language models to follow human instructions through supervised learning on instruction-output pairs.
-English IFT datasets include the FLAN collection(Wei et al., 2022), which scaled instruction tuning to over 1,800 tasks and chain-of-thought prompting, and Self-Instruct (Wang et al., 2023), which demonstrated effectiveness of LLM-generated synthetic data, leading to Stanford Alpaca (Taori et al., 2023) and the enhanced Alpaca-GPT4 (Peng et al., 2023).
-Subsequent work expanded through conversational approaches like Vicuna (Chiang et al., 2023) using ShareGPT conversations, WizardLM’s Xu et al.
-(2024) Evol-Instruct methodology, and LIMA’s (Zhou et al., 2023a) demonstration that 1K curated prompts suffice for effective tuning.
-The ORCA series (Mukherjee et al., 2023; Mitra et al., 2023) introduced explanation tuning and Prompt Erasure techniques, culminating in OrcaAgent-Instruct (Mitra et al., 2024a) with 25.8M synthetic instruction-response pairs.
-Multilingual Instruction Fine-tuning Datasets: Multilingual instruction-following capabilities have been developed through three primary methodologies.
-Translation-based methods create multilingual datasets by machine translation of existing English instruction data.
-Bactrian-X (Li et al., 2023) translates Alpaca (Taori et al., 2023) and Dolly (Conover et al., 2023) to produce 3.4M pairs across 52 languages, mAlpaca (Chen et al., 2024) demonstrates cost-effective multilingual tuning through Alpaca translation, and LIMA-X achieves 9.9% cross-lingual improvements using $\sim$1K instructions.
-Template-based approaches like M2Lingual leverage Evol-guided taxonomy (Xu et al., 2024) for synthetic generation across 70 languages.
-Hybrid strategies combine multiple techniques: Aya-Collection (Singh et al., 2024b) integrates crowdsourcing across 65 languages with templating and translation, repurposing xP3 (Muennighoff et al., 2023), Flan (Longpre et al., 2023), and Dolly datasets while employing NLLB 3.3B (Team et al., 2022) for n-way translation.
-Similarly, IndicAlign aggregates 74.7M prompt-response pairs across 20 Indian languages through aggregation of existing datasets, translations via the IndicTrans2 (Gala et al., 2023) model, synthetic conversation generation from India-centric Wikipedia, and crowdsourcing.
-Data Generation Strategies: Most prior efforts rely on output distillation from advanced LLMs like GPT-4, where models learn from teacher-generated responses.
-However, recent approaches have explored alternative strategies to address distillation limitations.
-Instruction Backtranslation (Li et al., 2024) starts with a model fine-tuned on seed data to generate instruction prompts for web documents, followed by self-curation to select high-quality examples.
-This leverages web content diversity while avoiding over-reliance on teacher models through iterative self-augmentation and self-curation.
-Instruction Back-and-Forth Translation (Nguyen et al., 2024) extends backtranslation by generating synthetic instructions and then rewriting responses using an LLM to improve quality based on initial documents.
-This combines web information diversity with high-quality response generation, outperforming distillation and yielding better win rates than ShareGPT and Alpaca-GPT4 (Peng et al., 2023).
-Limitations of prior multilingual IFT Datasets: Existing multilingual instruction datasets suffer from critical limitations that hinder effective cross-lingual alignment.
-Translation-based approaches like Bactrian-X, mAlpaca, and LIMA-X focus primarily on basic instruction-following while excluding advanced reasoning capabilities and lack demographic or cultural representation for local alignment.
-These datasets suffer from translation inaccuracies due to context-length limitations of traditional sentence-level translation methods (NLLB 3.3B and IndicTrans2) which fail to capture linguistic nuances and introduce errors that can potentially creep in during training.
-Despite impressive language coverage, Aya-Collection provides limited culturally-specific content, while IndicAlign sources a large proportion of data from WordNet (Miller, 1994) with a focus on question-answering offering limited task diversity.
-Most existing datasets are short-context and single-turn, failing to address real-world usage patterns including long-context or multi-turn dialogue capabilities.
-This can significantly constrain model performance in practical applications that demand sustained, contextually-aware interactions.
+English Instruction Fine-tuning (IFT).
+IFT adapts pre-trained LMs to follow instructions using instruction–response pairs (Ouyang et al., 2022).
+Early English IFT corpora include FLAN (Wei et al., 2022), which scaled to >>1.8k tasks with CoT prompting, and Self-Instruct (Wang et al., 2023), which showed the viability of LLM-generated synthetic data and inspired Stanford Alpaca (Taori et al., 2023) and Alpaca-GPT4 (Peng et al., 2023).
+Follow-on work explored conversational and curation-heavy routes: Vicuna (Chiang et al., 2023), WizardLM’s Evol-Instruct (Xu et al., 2024), LIMA (Zhou et al., 2023) showing that $\sim$1k high-quality prompts suffice, and the ORCA series (Mukherjee et al., 2023; Mitra et al., 2023), which introduced explanation tuning and prompt erasure, culminating in OrcaAgent-Instruct with 25.8M synthetic pairs (Mitra et al., 2024a).
+Multilingual IFT Datasets.
+Multilingual instruction-following has been pursued via (i) translation, (ii) template/synthesis, and (iii) hybrids.
+Translation-focused efforts include Bactrian-X (Li et al., 2023), which translates Alpaca (Taori et al., 2023) and Dolly (Conover et al., 2023) into 3.4M pairs over 52 languages, and mAlpaca (Chen et al., 2024), which translates Alpaca.
+Template-driven generation such as M2Lingual extends Evol-guided taxonomies (Xu et al., 2024) to 70 languages.
+Hybrid pipelines combine crowdsourcing, templating, and translation: Aya-Collection (Singh et al., 2024b) integrates crowd data across 65 languages with repurposed xP3 (Muennighoff et al., 2023), FLAN (Longpre et al., 2023), and Dolly, using NLLB 3.3B (Team et al., 2022) for N-way translation; IndicAlign aggregates 74.7M prompt–response pairs for 20 Indian languages via dataset aggregation, IndicTrans2-based translation (Gala et al., 2023), synthetic conversations from India-centric Wikipedia, and crowdsourcing.
+Data Generation Strategies.
+Most prior work distills outputs from stronger teachers (e.g., GPT-4).
+Recent alternatives mitigate distillation limits by leveraging diverse web content with self-augmentation and self-curation: Instruction Backtranslation (Li et al., 2024) synthesizes instructions from documents, and Back-and-Forth Translation (Nguyen et al., 2024) iteratively rewrites responses with LLMs, often outperforming pure distillation.
+Limitations of Prior Multilingual IFT.
+Translation-heavy datasets (e.g., Bactrian-X, mAlpaca, and even curation-focused LIMA-X) tend to emphasize basic instruction following, underrepresent advanced reasoning, and provide limited demographic/cultural grounding.
+Sentence-level MT (e.g., NLLB 3.3B, IndicTrans2) can introduce context-loss and subtle errors that propagate during training.
+Despite broad coverage, Aya-Collection contains comparatively little culturally specific content, while IndicAlign relies heavily on WordNet (Miller, 1994) and QA-style prompts, limiting task diversity.
+Finally, most corpora are short-context and single-turn, leaving long-context and multi-turn underexplored.
+Reasoning Subset (13 Indian Language & English) Generative Subset (13 Indian Language & English)
+Category Total Drop (%) Final Category Total Drop (%) Final
+Analytical R 350K 0.047 349.8K Logical R 229.4K 1.459 226.0K
+Brain Teaser 700K 0.043 699.7K Multihop QA 229.4K 1.459 226.0K
+Fermi 350K 0.015 349.9K Creative Writing 229.4K 1.459 226.0K
+Fs-CoT-Flow 350K 3.769 336.8K Multi-turn Dialogue 229.4K 1.611 225.7K
+Math 2.80M 0.035 2.80M Summarization 229.4K 1.526 225.9K
+MCQ 1.40M 0.135 1.40M Translation (to En) 229.4K 0.641 227.9K
+Reading Comp. 700K 0.379 Translation (from En) 229.4K 17.047 190.3K
+Text Classification 700K 1.878 686.9K Cultural MHR 375.7K 0.347 374.4K
+Causal R 229.4K 1.453 226.0K
+Total 7.35M 7.32M Total 2.21M 2.15M
+Table 1: Document filtering statistics aggregated over all 13 Indic languages and English.
+Totals and final counts are reported using K (thousands) and M (millions) notation.
+Per-language counts are uniform within each category and described in the text.
+MHR denotes Multi-Hop Reasoning, R denotes Reasoning.
 3 Data Generation
-The Updesh dataset comprises two complementary subsets addressing distinct aspects of multilingual instruction following: reasoning data and open-domain generative data.
+We synthesize Updesh, a dataset covering 13 Indic languages—Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Nepali, Odia, Punjabi, Tamil, Telugu, and Urdu.
+For each language, Updesh includes two complementary subsets targeting distinct facets of multilingual instruction following: reasoning and open-domain generation.
 This design recognizes that reasoning capabilities are largely language- and culture-agnostic, making translation-based approaches suitable for tasks like mathematical problem-solving and logical inference (Shaham et al., 2024).
+We summarize the key design considerations guiding the curation of Updesh in Appendix B.1 along with Figure 6.
 Existing high-quality reasoning datasets such as OrcaAgent-Instruct and OrcaMath, thus are valuable resources for multilingual adaptation.
 However, generative capabilities requiring cultural awareness, linguistic naturalness, and factual grounding in local contexts cannot be adequately addressed through translation due to inherent Western-centric biases and lack of cultural specificity in existing datasets (Yao et al., 2024).
-Therefore, our generative subset employs a grounded approach that ensures factuality through Wikipedia-based content, maintaining linguistic naturalness through native language generation, and preserves cultural adherence through systematic curation of India-specific cultural artifacts.
-Reasoning Data: Inspired by prior work (Ahuja et al., 2025; Khan et al., 2024), we translate eight subsets of the OrcaAgent-Instruct (Mitra et al., 2024a) and OrcaMath (Mitra et al., 2024b) datasets into 13 Indic languages.
-Specifically, we consider seven reasoning-related subsets from OrcaAgent-Instruct [^2] https://huggingface.co/datasets/microsoft/orca-agentinstruct-1M-v1 along with the Math subset from OrcaMath [^3] https://huggingface.co/datasets/microsoft/orca-math-word-problems-200k (Table 1).
+Therefore, our generative subset employs a grounded approach that ensures factuality through Wikipedia content, maintaining linguistic naturalness through native language generation, and preserves cultural adherence through systematic curation of India-specific cultural artifacts.
+Reasoning Data
+Inspired by prior work (Ahuja et al., 2025; Khan et al., 2024), we translate eight subsets of the OrcaAgent-Instruct (Mitra et al., 2024a) and OrcaMath (Mitra et al., 2024b) datasets into 13 Indic languages.
+Specifically, we consider seven reasoning-related subsets from OrcaAgent-Instruct [^1] https://huggingface.co/datasets/microsoft/orca-agentinstruct-1M-v1 along with the Math subset from OrcaMath [^2] https://huggingface.co/datasets/microsoft/orca-math-word-problems-200k (Table˜7).
 Both datasets have been attributed to induce significant chain-of-thought and reasoning capabilities in models during instruction-tuning without the need for specific preference optimization.
 We employ Llama-3.1-405B-Instruct for selective translation given its strong coverage in Indian languages and instruction-following capabilities that enable adaptation to various conversational styles (Sankar et al., 2025).
-Decoding is performed using nucleus sampling with top_p=0.95top\_p=0.95 and temperature set to $1.0$.
-After generation, all outputs undergo a series of quality control filters, described in detail below.
-Table 1: Reasoning task categories and their descriptions
-Task Type Description
-Analytical Reasoning MCQ-style questions requiring step-by-step logical inference
-Multiple-Choice Questions General-purpose problems across diverse knowledge domains
-Fermi (Guesstimation) Open-ended estimation problems using logical assumptions
-Few-Shot Chain-of-Thought Tasks with 4-5 in-context examples for learning
-Brain Teasers Puzzles stimulating lateral thinking and creativity
-Text Classification Categorization tasks for predefined labels
-Math Grade-school arithmetic, algebra, and geometry word problems
-Open-Domain Generative Data: Synthesizing generative data presents greater challenges than translating existing content, as it introduces increased risks of hallucinations, factual inaccuracies, and demographic misalignment.
-We conducted a comparative analysis between Llama-3.1-405B-Instruct and Qwen3-235B-A22B models across both reasoning and non-reasoning paradigms to evaluate qualitative differences in generated outputs.
-Manual assessment revealed that the Qwen3-235B-A22B model demonstrated superior proficiency in generative tasks and excelled in complex instruction following owing to its strong reasoning traces.
-Therefore, we used the Qwen3-235B-A22B for our generative data curation process.
-Drawing inspiration from instruction backtranslation techniques (Li et al., 2024), our methodology begins with unlabelled content to construct questions, followed by LLM-generated answers.
-To ensure generation diversity and contextual grounding that maintains factual accuracy and demographic relevance, we leverage Wikipedia pages in respective target languages as our knowledge base.
-Table 2 summarizes the eight generative task categories.
-As demonstrated, some tasks required two phases, which indicates two LLM inferences in order to generate the synthetic data instance.
-Table 2: Generative task categories with synthesis methods, phases, and model configuration
-Task Type Synthesis Method Phases Qwen3-Mode
-Logical Reasoning Generate implicit inferences from text passages (1) Direct inference generation Reasoning
-Multi-Hop QA Create questions requiring information synthesis across text segments (1) Question generation (2) Answer generation Reasoning
-Creative Writing Transform factual content into engaging narratives (1) Generate creative piece (2) Generate eliciting prompt Reasoning
-Multi-Turn Dialogue Agentic workflows with 3-5 turn conversations between personas (1) Generate dialog adhering to personas (2) Generate natural prompt Non-reasoning
-Summarization Generate summaries preserving key information across languages (1) Direct summary generation Non-reasoning
-Machine Translation Cross-lingual conversion maintaining cultural context (1) Direct translation Non-reasoning
-Causal Reasoning Identify and explain cause-effect relationships in text (1) Direct causal analysis Reasoning
-Further, to ensure cultural representation in synthetic data generation, we systematically curated culturally relevant content from Wikipedia using the MediaWiki API.
-Following (Yao et al., 2024)’s cultural taxonomy framework, we traversed Wikipedia category structures starting from Category:Culture of India and Category:Culture of India by state or union territory, exploring 2-3 levels deep.
-This yielded two complementary datasets: a general collection (54 categories, 19,143 artifacts) covering pan-Indian elements, and a state-wise collection (28 states + 8 union territories, 37,227 artifacts) with region-specific content spanning festivals, cuisine, traditional arts, architecture, and religious practices.
-We sampled 26.8K cultural artifacts from this collection to create multi-hop question-answer pairs for synthetic data generation.
-Data Filtering After generating approximately 7.5 million data points across 13 languages for both the Reasoning and Open-Domain subsets, manual validation was not feasible, therefore, following the approach of Shen et al.
+Post generation, all outputs undergo strict quality checks, as described in detail below.
+Figure 2: Human LLM-judge agreement across evaluation metrics, revealing differences across dimensions.
+Open-Domain Generative Data
+Synthesizing generative data poses greater challenges than translation due to increased risks of hallucinations, factual inaccuracies, and demographic misalignment.
+We compared Llama-3.1-405B-Instruct and Qwen3-235B-A22B across reasoning and non-reasoning paradigms, finding Qwen3-235B-A22B superior for generative tasks and complex instruction following due to stronger reasoning traces.
+(Also supported by Chiang et al.
+(2024)).
+Inspired by instruction backtranslation techniques (Li et al., 2024), we construct questions from unlabelled content followed by LLM-generated answers.
+To ensure diversity, contextual grounding, factual accuracy, and demographic relevance, we leverage Wikipedia pages in target languages as our knowledge base.
+Table 8 summarizes eight generative task categories, with some requiring two LLM inference phases.
+Further, to ensure cultural representation, we systematically curated culturally relevant content from Wikipedia using the MediaWiki API.
+Following (Yao et al., 2024)’s cultural taxonomy, we traversed Wikipedia categories from Category:Culture of India and Category:Culture of India by state or union territory, exploring 2-3 levels deep.
+This yielded diverse region-specific content spanning festivals, cuisine, traditional arts, architecture, and religious practices.
+We sampled 26.8K cultural artifacts to create multi-hop question-answer pairs for synthetic data generation.
+For English, we reuse the reasoning data from OrcaAgent-Instruct (Mitra et al., 2024a) and OrcaMath (Mitra et al., 2024b) as-is, while generating the English generative subset from scratch using the same pipeline as for the Indian languages.
+Data Filtering
+After generating data points at scale across 13 languages for both the Reasoning and Open-Domain subsets, manual validation was not feasible, therefore, following the approach of Shen et al.
 (2025), we employed automated quality checks but use the standard threshold-based method instead of their anomaly detection-based method.
@@
 For the English-to-XX translation tasks, Assamese had the highest drop rate as the model frequently generated Bengali text instead, likely due to the shared script and similarity between these languages, and because Assamese is a low-resource language.
-Table 3: Document filtering statistics averaged over all Indic Languages
-Reasoning Subset Generative Subset
-Category Total Drop Rate (%) Category Total Drop Rate (%)
-analytical_reasoning 25,000 0.047 logical_reasoning 16,384 1.459
-brain_teaser 50,000 0.043 multihop_QA 16,384 1.459
-fermi 25,000 0.015 creative_writing 16,384 1.459
-fs_cot_flow 25,000 3.769 Multi_turn_Dialog 16,384 1.611
-math 200,035 0.035 summarization 16,384 1.526
-mcq 99,986 0.135 machine_translation_en 16,384 0.641
-rc 50,000 0.379 machine_translation_xx 16,384 17.047
-text_classification 50,000 1.878 cultural_multihop_reasoning 26,834 0.347
-causal_reasoning 16,384 1.453
-5 Data Quality Analysis
-5.1 Translation Quality Assessment for the Reasoning Subset
-For the reasoning subset, we conducted large-scale selective translation utilizing the Llama-3.1-405B-Instruct.
-Given that the input often consists of long-context and non-standard text, it is crucial to rigorously evaluate the quality of the generated translations.
-To this end, following established methodologies the generated output in the target language is translated back into English and compared with the original English source.
-Due to the substantial volume of data, we randomly sample 4,096 instances from each subset for every language to ensure a representative evaluation.
-For the backtranslation step, we employed the more efficient Llama-3.3-70B-Instruct, which offers faster inference and serves as a conservative lower bound on translation quality.
-We assessed translation fidelity using standard metrics such as ChrF (Popović, 2015), implemented using the SACREBLEU library (Post, 2018) for consistency.
-Table 9 in the Appendix demonstrates consistently high backtranslation scores across all languages and subsets, indicating robust translation quality.
-5.2 Data Quality Assessment for Generative Data
-Although large language models (LLMs) have emerged as scalable evaluators following the popular LLM-as-a-judge paradigm, their reliability in culturally nuanced and low-resource language settings still remains limited  (Watts et al., 2024).
-Our study combined LLM-based assessment with native-speaker human annotation to quantify data quality and establish inter-annotator agreement.
-We employed a stratified sampling strategy to select 100 data points per task category: Creative Writing, Cultural Multi-Hop Reasoning, Multi-Turn Dialogue, and aggregated reasoning tasks (Logical, Causal, and Multi-Hop Reasoning) across five Indian languages (Assamese, Gujarati, Hindi, Malayalam and Punjabi) resulting in a total of 2000 data points.
-Our sampling methodology ensures representative coverage by partitioning responses into quintiles based on the output length and maintaining proportional sampling across buckets to preserve the original length distribution while capturing variance across responses.
-Human evaluations were conducted by native speakers through an external agency (Table 7).
-We employed GPT-4o as an automated evaluator using identical evaluation protocols to allow direct comparison.
-We devised the following metrics to capture multidimensional quality aspects:
-  • Creative Writing: Instruction adherence, linguistic fluency, narrative coherence
-  • Reasoning Tasks: Answer adequacy, contextual adherence, instruction compliance, fluency and readability, problematic content detection, cultural relevance assessment (cultural multi-hop only)
-  • Multi-Turn Dialogue: Persona consistency, topical coherence, linguistic plausibility, repetition detection, toxicity screening, instruction adherence
-  • Summarization: Content coverage, factual accuracy, conciseness, coherence and logical flow, stylistic appropriateness
-  • Translation: Semantic correctness, fluency corectness, domain appropriateness, style and tone, completeness
-Further details on the metrics and their definitions could be found in the Appendix in Table LABEL:tab:rubrics.
-All metrics utilize a three-point Likert scale (0-2) with detailed rubrics provided to both human and automated evaluators to ensure assessment consistency.
-The prompts used for the LLM evaluations could be found in the Appendix A.7.
-The guidelines given to the human annotators have been added in the supplementary material.
-Quantitative analysis of human evaluations reveals very high data quality: across 10,000 individual metric assessments, human evaluators assigned only 27 scores of zero (0.27%), indicating that the overwhelming majority of generated content meets acceptable quality thresholds.
-We plan to publicly release all evaluation data, including both human and GPT assessments, to ensure complete transparency and support future research on calibration of LLM-based evaluators.
-Figure 1: Human LLM-judge agreement across evaluation metrics, revealing differences across dimensions.
-Inter-Annotator Agreement: To quantify the reliability of automated evaluation, we computed the percentage agreement between the LLM-judge and the human evaluators.
-We find significant variance in agreement across metrics (Figure 1).
-Agreement substantially deteriorates for culturally and linguistically nuanced assessments, particularly for linguistic plausibility evaluation and in repetition detection in long dialogue sequences.
-Conversely, well-defined objective criteria, such as toxicity detection and problematic content identification, demonstrate robust inter-annotator alignment.
-Since our data is generated using benign prompts and topics, toxic or problematic content is expected to be very rare or nonexistent.
-These results are consistent with earlier research on the challenges of subjective versus objective evaluation and reveal shortcomings in current LLM-judges for assessing culturally sensitive content  (Watts et al., 2024).
-The distribution of scores across tasks and languages for both humans and LLM can be found in Appendix  A.2.
+4 Dataset Quality Analysis (Q-A)
+4.1 Q-A for Reasoning Data
+For the reasoning subset, we performed large-scale selective translation using Llama-3.1-405B-Instruct.
+Given inputs with long contexts and non-standard text, we rigorously evaluated translation quality through backtranslation.
+We randomly selected 4,096 samples per subset and language, backtranslated them to English using Llama-3.3-70B-Instruct (chosen for faster inference and conservative quality bounds), and compared with original sources.
+Translation fidelity was measured using ChrF (Popović, 2015) via SACREBLEU (Post, 2018).
+Table 10 shows consistently high backtranslation scores across all languages, confirming robust translation quality.
+4.2 Q-A for Generative Data
+While large language models (LLMs) have become scalable evaluators under the LLM-as-a-judge paradigm, their reliability in culturally nuanced and low-resource settings remains limited (Watts et al., 2024; Whitehouse et al., 2025).
+We therefore combined LLM evaluation with native-speaker annotation and measured inter-annotator agreement.
+Using stratified sampling, we drew 100 instances per category—Creative Writing, Cultural Multi-Hop Reasoning, Multi-Turn Dialogue, and aggregated reasoning (Logical, Causal, MultiHop)—across five languages (Assamese, Gujarati, Hindi, Malayalam, Punjabi), yielding 2K items.
+Sampling preserved response-length distributions via quintile bucketing.
+Native-speaker evaluations were conducted via an external agency (Table 6), while GPT-4o served as an automated evaluator using identical protocols for comparability.
+We defined task-specific, multidimensional metrics using a 3-point Likert scale (0–2) with consistent rubrics (Section B.3), ensuring a thorough quality check of the generated data.
+Prompts and annotation guidelines are provided in Appendix C.1 and supplementary material.
+Across 10K individual ratings, only 27 received a zero (0.27%), indicating uniformly high data quality.
+We will publicly release all evaluation data—including both human and GPT assessments—to promote research on calibrating LLM-based evaluators.
+Inter-Annotator Agreement To assess the reliability of automated evaluation, we computed percentage agreement between the LLM-judge and human evaluators.
+Agreement varies notably across metrics (Figure 2), declining for culturally and linguistically nuanced aspects such as linguistic plausibility and repetition detection in long dialogue sequences.
+In contrast, objective criteria like toxicity detection and problematic content identification show strong alignment.
+As the data is derived from benign prompts and topics, toxic or problematic instances are expected to be rare or absent.
+These findings align with prior work on the difficulty of subjective versus objective evaluation and highlight persistent limitations of current LLM-judges in assessing culturally sensitive content (Watts et al., 2024).
+The complete distribution of scores across tasks and languages for both evaluations is provided in Appendix B.4.
+Figure 3: Evaluation plots for models finetuned on Updesh vs existing datasets
+Figure 4: Model Performance Landscape: NLU vs.
+NLG vs.
+Win Counts.
+The horizontal axis represents the average NLU (accuracy between 0-100), while the vertical axis represents the average NLG score (ChrF between 0-100).
+The size of each bubble corresponds to the number of specific datasets (12 tasks evaluated) where that the model outperformed all others.
+Updesh model (green) demonstrates the most dominant position with high generation scores and the largest number of task wins across both Llama and Phi settings.
+Figure 5: NLG performance across 16 out-of-domain Indic languages on Flores.
+Updesh (red) achieves the highest average scores on both Llama-3-8B and Phi-4 architectures, outperforming standard baselines (Zero-shot) and comparable instruction-tuned models (Bactrian, Aya, IndicAlign).
 5 Downstream Tasks Evaluation
@@
 We used the Axolotl framework [^3] https://github.com/axolotl-ai-cloud/axolotl for all fine-tuning runs.
-Details regarding the hyperparameters and compute resources used can be found in Appendix A.4.
-Baselines: We fine-tuned both Llama-3.1-8B and Phi4-14B on three high-quality, open-source IFT datasets that collectively cover nearly all the Indic languages included in our work: the Aya-Collection (Singh et al., 2024b), IndicAlign (Khan et al., 2024) and the Bactrian-X dataset (Li et al., 2023).
-To the best of our knowledge, these are the only open datasets that offer both broad language coverage and instruction-following format, and there is a lack of other large-scale open IFT datasets for Indian languages.
-The Bactrian-X dataset provides training samples for 10 out of our 13 target languages, excluding Assamese, Kannada and Punjabi.
-The Aya-Collection dataset includes all target languages except Punjabi.
-Since Aya-Collection contains millions of samples per language, we uniformly sub-sampled it to construct a balanced dataset of approximately 7 million samples which is comparable in scale to Updesh.
-For IndicAlign, the WordNet subset ($\sim$97M pairs) is disproportionately large, less diverse, and redundant, so we downsampled it to one instance per entry, yielding about 7.3M training pairs while retaining all other subsets.
-Downstream Tasks: Our evaluation framework consists of three task categories to comprehensively assess model capabilities.
+Details regarding the hyperparameters and compute resources used can be found in Appendix B.6.
+Baselines
+We fine-tuned both Llama-3.1-8B and Phi4-14B on three high-quality, open-source IFT datasets: Aya-Collection (Singh et al., 2024b), IndicAlign (Khan et al., 2024) and Bactrian-X (Li et al., 2023).
+To the best of our knowledge, these are the only open datasets that offer both broad language coverage and an instruction-following format.
+Bactrian-X covers 10 of our 13 languages (excluding Assamese, Kannada, Punjabi), while Aya-Collection includes all except Punjabi.
+Since Aya-Collection contains millions of samples per language, we uniformly sub-sampled it to 7M samples to create a balanced dataset comparable to Updesh.
+Similarly, for IndicAlign, since the WordNet subset ($\sim$97M pairs) is disproportionately large, less diverse, and redundant, we downsampled it to one instance per entry, yielding 7.3M training pairs when combined with its remaining subsets.
+Downstream Tasks
+Our evaluation framework consists of three task categories to comprehensively assess model capabilities.
 Natural language understanding (NLU) tasks use multiple-choice questions to measure comprehension and reasoning through likelihood-based scoring.
-Natural language generation (NLG) tasks like translation and summarization, assess models’ ability to generate coherent and contextually appropriate outputs.
-For instruction-following evaluation, we translate established benchmarks (IFEval, IFBench) using GPT-4o (prompts in A.6), enabling assessment of both cross-lingual instruction-following (English responses to non-English instructions) and native language instruction-following.
-This design systematically identifies model strengths and weaknesses across diverse tasks, providing a holistic performance assessment.
-Dataset details are in Appendix A.5.
+Natural language generation (NLG) tasks, such as translation and summarization, assess models’ ability to generate coherent and contextually appropriate outputs.
+We augment standard dataset-NLU and NLG evaluations with comparative evaluations to understand model win rates.
+This design identifies model strengths and weaknesses across diverse tasks, providing a holistic performance assessment.
+Dataset details are in Table 2.
+Dataset Source
+NLU MMLU Indic (MMLU-I) SarvamAI (2025)
+ARC Indic (ARC-I) SarvamAI (2025)
+BoolQ Indic (BoolQ-I) SarvamAI (2025)
+TriviaQA Indic (TVQA-I) SarvamAI (2025)
+BeleBele (Bele) Bandarkar et al. (2024)
+INCLUDE (INCL) Romanou et al. (2025)
+Global MMLU (GMMLU) Singh et al. (2025)
+NLG Extreme Summarization (Xsum) Singh et al. (2024a)
+Flores English to Others (Flores EnXX) Goyal et al. (2022)
+Flores Others to English (Flores XXEn)
+IN22-Conv (IN22-Conv-Doc) - EnXX Gala et al. (2023) Gumma et al. (2025)
+IN22-Conv (IN22-Conv-Doc) - XXEn
+Table 2: Evaluation datasets
 5.1 Results
-Table 4: NLU task performance. All entries are accuracy (higher is better).
-Model Setting MMLU MMLU-I MILU ARC-I BoolQ-I TVQA-I Bele INCL GMMLU
-Llama Base 0.23 0.23 0.26 0.25 0.62 0.27 0.24 0.25 0.23
-Bactrian 0.63 0.37 0.42 0.26 0.65 0.66 0.41 0.39 0.42
-Indic-Align 0.60 0.34 0.42 0.27 0.71 0.59 0.32 0.34 0.39
-Aya 0.59 0.33 0.39 0.28 0.76 0.51 0.47 0.34 0.37
-Updesh 0.51 0.34 0.37 0.30 0.78 0.46 0.51 0.31 0.37
-Phi4 Base 0.78 0.43 0.52 0.30 0.80 0.68 0.55 0.45 0.55
-Bactrian 0.78 0.40 0.48 0.29 0.74 0.67 0.48 0.42 0.48
-Indic-Align 0.77 0.45 0.51 0.32 0.76 0.62 0.52 0.45 0.54
-Aya 0.78 0.45 0.53 0.33 0.73 0.72 0.61 0.48 0.53
-Updesh 0.78 0.49 0.54 0.32 0.81 0.60 0.74 0.47 0.57
-Performance on NLU Tasks: Table 4 shows the performance of both base models fine-tuned on all three IFT datasets, along with the performance of the base models on NLU tasks.
-Phi4-Updesh stands out as the most consistently strong configuration, attaining the best overall scores on MMLU-I, MILU, BoolQ-I, BeleBele, INCL, and GlobalMMLU.
-For the rest of the tasks, different models excel - in case of the Llama base model, Bactrian-X leads knowledge-intensive tasks, attaining the highest on MMLU and TVQA-I, while Updesh excels in instruction-style reasoning benchmarks (ARC-I, BoolQ-I) and performs best on reading comprehension tasks such as BeleBele.
-Phi4’s MMLU performance appears saturated across all fine-tuning baselines, suggesting a ceiling effect from pre-training knowledge or lack of coverage within all the training mixtures.
-Overall, these results demonstrate that different IFT datasets offer distinct advantages across various NLU capabilities.
-Table 5: NLG task performance. All entries are ChrF scores (higher is better).
-Model Setting XSum Flores En-XX Flores XX-En IN22-Conv-Doc
-Llama Base 0.16 1.00 36.91 0.45
-Bactrian 0.21 23.69 46.31 14.09
-Indic-Align 9.26 25.18 2.29 10.98
-Aya 0.24 22.60 0.53 6.73
-Updesh 17.39 32.63 46.84 21.25
-Phi4 Base 11.32 23.03 51.39 22.80
-Bactrian 0.31 25.93 47.75 15.45
-Indic-Align 0.27 27.48 1.77 13.82
-Aya 0.37 26.51 7.63 14.01
-Updesh 10.21 34.58 53.29 18.74
-Performance on NLG Tasks: Table 5 demonstrates that Updesh is clearly superior to the other datasets on NLG tasks.
-Llama-Updesh emerges as the best configuration, achieving the highest performance across all generative tasks with substantial margins over all competing datasets as well as the base model.
-In case of Phi4, Updesh delivers the strongest translation performance, building effectively on the already capable base model.
-However, summarization results are more nuanced: the base model leads on XSum with Updesh being competitive, while the other fine-tuned models perform poorly.
-On long-context conversational translation (IN22-Conv-Doc) outperforms all fine-tuned variants, though it slightly lags behind the Phi4 Base model.
-Overall, we observe that the long-context generative training data in Updesh provides significant benefits for generative tasks.
-Table 6: Instruction-following capabilities. All entries are accuracy (higher is better).
-Model Setting mIFEval mIFBench
-Native in English Native in English
-Prompt Inst. Prompt Inst. Prompt Inst. Prompt Inst.
-Llama Base 0.24 0.36 0.18 0.31 0.19 0.21 0.16 0.17
-Bactrian 0.17 0.27 0.14 0.25 0.17 0.18 0.11 0.13
-Indic-Align 0.16 0.24 0.14 0.23 0.17 0.19 0.16 0.17
-Aya 0.13 0.22 0.11 0.21 0.18 0.19 0.18 0.19
-Updesh 0.20 0.30 0.18 0.28 0.16 0.19 0.14 0.15
-Phi4 Base 0.22 0.32 0.19 0.30 0.14 0.16 0.13 0.14
-Bactrian 0.22 0.32 0.19 0.30 0.20 0.21 0.15 0.16
-Indic-Align 0.16 0.25 0.15 0.24 0.15 0.17 0.14 0.16
-Aya 0.15 0.25 0.14 0.24 0.20 0.21 0.19 0.20
-Updesh 0.22 0.32 0.19 0.29 0.19 0.22 0.16 0.18
-Performance on Instruction Following Tasks: Table 6 presents results on the GPT-4o translated multilingual versions of IFEval and IFBench, which evaluate fine-grained constraint adherence through highly specific requirements like character positioning, word counts, and formatting specifications - constraints notably absent in any training datasets.
-Native refers to prompts as well as expected output in regional language, while in English uses regional language prompts but expects English responses.
-Prompt-level evaluation measures whether all instructions in a prompt are followed, while instruction-level evaluation measures individual instruction adherence.
-Updesh emphasizes general instruction following involving task completion and output formatting, making these benchmarks an effective OOD evaluation for precise constraint adherence.
-For Llama, fine-tuning generally degrades performance, but Updesh shows smaller drops (mIFEval: 3-4 point lesser degradation than next best - Bactrian; mIFBench: comparable or better performance than other baselines, lags Aya on in English tasks).
-Phi4 exhibits different dynamics - Updesh achieves positive gains on mIFBench over the Base model (+4/+2 points Native; +1/+3 points English) while maintaining base-level performance on mIFEval, outperforming other fine-tuning approaches (Aya, Indic-Align) that show mixed or negative changes.
-Updesh provides robust instruction-following with minimal catastrophic forgetting compared to other baselines.
-Language-wise trends: Figure 2 shows a clear trend: performance improves as language resources increase for both models and task types (NLU accuracy and NLG ChrF).
-We illustrate this with the resource taxonomy Joshi et al.
-(2020), ordered from 0 (lowest-resource) to 5 (high resource).
-The relative improvements with Updesh are most pronounced in low/mid-resource bands, effectively narrowing the gap to high-resource languages.
-The base model (without fine-tuning) is the weakest, with Bactrian-X and IndicAlign often competitive in the medium resource languages while Aya-Collection is more variable.
-Notably, generation benefits more from data quality: NLG improvements increase more dramatically as languages become lower-resourced, most clearly for Phi4-Updesh.
-Across language-resource tiers, Updesh is consistently strongest, especially at higher resource levels, reaching about $\sim$0.65 for Llama (NLU), $\sim$0.76 for Phi-4 (NLU) in terms of average accuracy, and the highest ChrF for Phi-4 (NLG).
-Figure 2: NLU and NLG task performance grouped by language–resource class taxonomy from Joshi et al. (2020). Updesh yields the largest relative gains in low/mid-resource languages
-6.2 Discussion
-Quality evaluation: Our comprehensive quality evaluation ranges from basic filters to nuanced human evaluation.
-Experiments with LLM-based evaluators show that while LLM-judges do well on some metrics, they struggle with subtle multilingual and multicultural contexts.
-For example, they often misjudge how a language is typically spoken or written, and tend to lost track of persona consistency in longer dialogues.
-These issues motivate the need for scalable and reliable evaluation methods that go beyond surface-level checks.
-Task-type sensitivity (NLU vs.NLG): We observe variation by task type - for instance, on NLU-style tasks (usually multiple-choice), no configuration is a clear winner across all datasets; performance depends on the base model and the specific benchmark.
-In contrast, on NLG-style tasks that score free-form generations, models trained on Updesh consistently perform strongly.
-This pattern suggests that the quality control and discourse control learned during training is transferred more directly to NLG than to choice-based NLU.
-Evaluation protocol differences matter: Most NLU evaluations compute correctness via likelihood comparisons over (context, option) pairs, effectively testing discrimination among short alternatives.
-NLG evaluations, by design, require longer generations and measure faithfulness, adequacy, and fluency under open-ended prompting.
-The divergence in scoring objectives partially explains why a dataset that emphasizes long-form instruction-following (like Updesh) yields larger gains on NLG than on likelihood-based NLU.
-We find that the Bactrian dataset tends to perform best on NLU-style benchmarks, which can be attributed to format alignment and a larger proportion of MCQ or short-from prompts.
-Distribution mismatches between train and test: Updesh contains longer-context samples and lacks short-text MCQ-style items; its task composition emphasizes multi-step, generative instruction-following.
-Even the MCQ subset which is part of the reasoning data in Updesh consists of long-form answers to MCQs, with close to no representation of short-text or direct answers.
-In contrast, many NLU benchmarks, however, are MCQ-heavy and short-context.
-This composition shift (task format, length, and response style) likely reduces the relative NLU gains for Updesh-trained models while amplifying NLG gains, showing that models generalize best to downstream tasks with evaluation formats that resemble their training distribution.
+Figure 3 presents a comparative performance analysis of the Llama and Phi-4 architectures across NLG and NLU tasks for a diverse set of 13 Indic languages.
+Broadly, we observe that models fine-tuned on the Updesh dataset (represented in green) consistently outperform existing baselines, including Bactrian-X, IndicAlign, and Aya-Collection.
+This performance advantage is particularly pronounced in NLG settings, where Updesh fine-tuned models demonstrate substantially higher performance across both high-resource languages like Hindi and Bengali, as well as lower-resource ones such as Assamese and Odia.
+Detailed results for the NLG tasks could be found in Table 12.
+In NLU tasks, Updesh maintains a competitive edge, often surpassing the strongest baselines highlighting the efficacy of the dataset in fostering robust multilingual understanding and generation capabilities.
+Figure 4 illustrates the comparative performance of the models across three distinct dimensions: Understanding (NLU), Generation (NLG), and overall robustness (Win Counts).
+While the NLU scores (x-axis) show a competitive landscape with tight clustering among fine-tuned models, the Updesh setting (green bubble) distinguishes itself significantly in generation tasks, consistently achieving the highest placement in terms of scores.
+Crucially, the bubble size indicates that Updesh secures the highest number of ‘wins’ - 7 for Llama3 and 8 for Phi-4 -far surpassing other baselines like Bactrian and Aya.
+We observe that Updesh has more pronounced NLU performance gains in the Phi4 setting compared to the Llama settings.
+Language-wise breakdowns and dataset-level averages are reported in Appendix B.8 (Tables 11 and 12).
+Evaluation on Unseen Languages
+To further understand robustness and cross-lingual transfer it is essential to test cross-lingual transfer to languages not seen in training dataset.
+In one of our NLG evaluation dataset(Flores) there are 16 languages which are not present in training data - Awadhi, Bhojpuri, Bodo, Chhattisgarhi, Garhwali, Haryanvi, Konkani, Maithili, Malvi, Manipuri, Marwari, Pashto, Rajasthani, Sanskrit, Santali and Tibetan.
+As illustrated in Figure 5, Updesh consistently outperforms all the other baselines across different model architectures.
+On Llama-3-8B, it achieves a strong NLG score of 24.93, clearly surpassing all counterparts, and the trend holds for Phi-4, where Updesh (24.10) again leads across baselines.
+These results demonstrate that Updesh ’s curation strategy enables robust cross-lingual transfer, even to languages unseen during training.
+Exact results on all languages could be found in Table 15.
+Setting NLU
+Avg NLG
+Avg Dataset
+Wins
+0-shot Baseline 28.64 12.69 0
+Updesh - R + G 43.97 32.28 6
+Updesh - G 48.48 19.73 7
+$\Delta$ (vs R+G) 4.42 -12.54
+Updesh - R 41.53 8.18 0
+$\Delta$ (vs R+G) -2.43 -24.10
+Llama
+Setting NLU
+Avg NLG
+Avg Dataset
+Wins
+0-shot Baseline 56.21 28.58 3
+Updesh - R + G 59.20 32.69 7
+Updesh - G 59.18 32.04 3
+$\Delta$ (vs R+G) -0.02 -0.64
+Updesh - R 56.79 1.80 0
+$\Delta$ (vs R+G) -2.41 -30.86
+Phi-4
+Table 3: Comparative performance of Updesh ablations.
+R and G denote the Reasoning and Generation subsets of the dataset, respectively.
+$\Delta$ values represent the performance difference compared to the full (R + G) model.
+Ablations on dataset composition
+To disentangle the relative contributions of the reasoning and generative subsets introduced in Sections 3, we conduct controlled ablations on the Updesh dataset.
+Specifically, we isolate the Reasoning (R) component - comprising translated subsets of the OrcaAgent-Instruct and OrcaMath datasets focused on multi-step reasoning and chain-of-thought supervision - and the Generative (G) component, which contains open-domain instruction-following and culturally grounded synthesis tasks derived from Indic Wikipedia content.
+We train individual models on each subset (Updesh-R and Updesh-G) and compare them against the full combined dataset (Updesh-R+G) to quantify their respective effects on NLU and NLG performance.
+Results presented in Table 3 reveal a clear interaction between dataset composition and model behaviour.
+For both Llama and Phi-4, isolating the generative subset slightly improves NLU (e.g., +3.78+3.78 for Llama) but substantially reduces NLG quality, suggesting that reasoning data - being of translated nature is beneficial for translation performance.
+Conversely, training exclusively on reasoning data leads to a marked decline in both NLU and NLG metrics (37.06-37.06 for Phi-4 NLG).
+These results demonstrate that our bottom-up data generation approach is superior to naive translation.
+Setting NLU
+Avg NLG
+Avg Dataset
+Wins
+0-shot Baseline 53.17 34.42 3
+Updesh - R + G 56.61 38.73 3
+Updesh - R + G (Seq len = 32K) 55.85 40.59 7
+$\Delta$ (vs R+G) -0.75 1.86
+Table 4: Ablations on training sequence length
+Ablations on training sequence length
+Additionally, given that many benchmark tasks contain shorter contexts, it is necessary to determine whether the long-context nature of the training data might have interference with short-context capabilities.
+Therefore, we evaluate a variant of the Updesh trained with a 32K444halved from the original 64K context window to examine the sensitivity of performance to sequence length constraints.
+Notably, we observe in 3, that the Phi-4 32K variant achieves the most balanced profile, securing the highest number of dataset wins ($7$) by maintaining strong NLU performance while further boosting NLG capabilities.
+This indicates that training sequence length is an important design consideration.
+Rank Model ELO Wins Losses Ties Win Rate
+1 Updesh-32K 1695.55 18760 6897 742 0.711
+2 IndicAlign 1658.50 17447 7991 657 0.669
+3 Updesh 1606.52 16168 8868 1111 0.618
+4 Updesh Reasoning 1537.19 12237 12371 1792 0.464
+5 Updesh Generative 1483.15 13752 11206 1442 0.521
+6 Bactrian-X 1311.22 6862 17806 1554 0.262
+7 aya 1207.86 2313 22400 1588 0.088
+Table 5: Overall ELO Rankings comparing Updesh variants against baseline models for Phi-4.
+The Updesh-32K model demonstrates superior performance, outperforming both internal variants and external baselines like Bactrian-X and Aya-Collection
+5.2 Comparative Cultural evaluations (ELO Rankings)
+In Section 5, we evaluate the trained models and baselines on well-established academic benchmarks, which primarily measure performance on standardized tasks.
+However, such evaluations do not fully capture how useful these models are for real-world user queries spanning diverse domains, nor do they adequately reflect their helpfulness to everyday users in culturally grounded scenarios.
+To address this gap, it is essential to perform robust comparative evaluations in nuanced cultural contexts, focusing on non-academic, real-world questions posed by users.
+[^5] Due to the high evaluation cost, we perform this analysis only on Phi-4 checkpoints; we focus on Phi-4 since it consistently outperforms Llama in our other evaluations.
+Thereby, we collect data following the collection process for Samiksha Hamna et al.
+(2025) and created a set of questions to assess the cultural relevance and helpfulness of LLM responses on practical, community-driven queries.
+Following the LLM-as-a-judge framework, we utilized GPT-4o to determine the superior response in pairwise comparisons between model checkpoints.
+Our evaluation encompasses 91,982 battles across seven models.
+To ensure statistical robustness, we evaluated all possible model pairings and randomized answer positions to mitigate positional bias, and calculate ELO ratings following the work in Boubdir et al.
+(2023).
+ELO ratings are calculated as the battles progress, where a higher rating indicates better comparative performance.
+This evaluation indicates that Updesh-32K significantly outperforms other baselines, with only IndicAlign showing competitive performance.
+Table 5 presents a comparative evaluation of model performance based on ELO ratings.
+The Updesh-32K model achieves the highest rating of 1696, establishing itself as the top-performing model in this evaluation set.
+It marginally outperforms Indic Align Cleaned (1659) and maintains a significant lead over the standard Updesh (1607).
+Notably, the Updesh models demonstrate substantial improvements over existing multilingual baselines, with the lead model scoring over 300 points higher than Bactrian-X Indic (1311) and nearly 500 points higher than the Aya-Collection Indic Sampled (1208), highlighting the efficacy of the Updesh dataset in providing more useful and grounded answers to India-centric, domain-specific queries.
+Alignment between ELO Scores and Automated Benchmarks
+As illustrated in Figure 11, there is a clear positive correlation between the ELO scores and the evaluated metrics across the board.
+Notably, the Updesh-32K (U32K) variant demonstrates superior performance, consistently clustering in the upper-right quadrant of all three plots.
+It achieves the highest ELO score ($\approx$ 1700) while simultaneously maintaining leading scores in NLU Average ($\approx$ 0.55), NLG Average ($\approx$ 35), and the total number of Dataset Wins.
+In contrast, baseline models such as Bactrian and Aya show mixed results; while Aya remains competitive in NLG tasks, it lags significantly behind U32K in the aggregate ELO ranking.
 6 Conclusion
@@
 We built Updesh, a 9.5M 13-language IFT dataset using a culturally grounded, bottom-up pipeline.
-Our comprehensive evaluation spanning data generation, quality assessment (human and LLM-as-judge), and downstream tasks, revealed that synthetic data can potentially bridge resource gaps.
-However, no single recipe dominates across tasks and models, motivating approaches tailored to diverse linguistic and cultural contexts.
-We will release the Updesh dataset, evaluation protocols, and detailed analyses to enable transparent, reproducible progress.
+Our comprehensive evaluation spanning data generation, quality assessment (human, LLM-as-judge), and downstream tasks, revealed that synthetic data can potentially bridge resource gaps.
+Results show Updesh dominates across tasks and models, for both NLU and NLG settings.
+We will release the Updesh dataset, evaluation protocols, and detailed analyses to enable future research.
+7 Limitations
+Lack of Reliable Data Quality Estimation for multilingual synthetic data
+Our comprehensive evaluation revealed that current LLM-based evaluators demonstrate variable reliability across quality dimensions, showing strong agreement with human judgments on objective metrics like toxicity detection (96-98%) but significantly lower concordance on nuanced aspects like fluency assessment and persona consistency (45-60%).
+This necessitates exercising caution when relying solely on LLM-based evaluations for quality estimation of multilingual synthetic data and highlights the need for more calibrated evaluators and robust evaluation frameworks.
+Cultural Authenticity and Knowledge Base Limitations
+Cultural authenticity remains challenging due to reliance on Wikipedia as the primary knowledge base, as many cultural customs and contextual nuances specific to under-represented Indian communities may lack sufficient documentation on Wikipedia, potentially resulting in incomplete cultural representations that might favor well-documented urban practices over rural or minority contexts.
+Lack of Specialized Benchmarking
+Limited benchmarks exist for evaluating cultural aspects and long-context/multi-turn capabilities in Indic languages, making systematic assessment of these crucial aspects difficult despite Updesh’s emphasis on these capabilities.
+While our framework covers general NLU, NLG specialized benchmarks for cultural knowledge and reasoning are needed to systematically evaluate and make progress.
+Although we conduct an comparative evaluation on real-world queries asked by human users, however we have used LLM evaluator for rating, while human evaluation would be a more fine-grained indicator of quality which we leave for future work.
 8 Ethical Considerations
-Our discussion of ethical considerations is guided by the framework proposed by Bender & Friedman (2018).
+Our discussion of ethical considerations is guided by the framework proposed by Bender and Friedman (2018).
 Institutional Process and Oversight
@@
 Table 6 summarizes the annotator demographics (education, region, age distribution, and gender).
-Each data worker was compensated at the rate of \$2 per data point.
+Each data worker was compensated at the rate of \$2 per data point, which is significantly higher than the average for an regular annotation task.
 Category Summary
@@
 Male: 4
-Table 7: Participant demographics summary
-Appendix A Appendix
-A.1 Rubrics used for the quality evaluation of the synthetic data
-Table 8: Evaluation Rubrics for AI Tasks
-Task Rubric Description
-Creative Writing Instruction adherence Assesses if the output strictly follows all constraints and guidelines provided in the prompt.
-Fluency Evaluates the naturalness, grammatical correctness, and readability of the generated text.
-Narrative coherence Checks for logical consistency in the plot, character development, and thematic elements.
-Reasoning Tasks Answer adequacy Determines if the final answer is correct, complete, and directly addresses the core question.
-Context adherence Measures whether the reasoning remains faithful to the provided context, avoiding external facts.
-Instruction adherence Verifies that the output’s structure, format, and steps match the user’s instructions.
-Fluency and readability Assesses the clarity, logical flow, and ease of understanding of the explanation.
-Problematic content and cultural relevance Scrutinizes the response for harmful stereotypes and ensures it is culturally sensitive and appropriate.
-Multi-turn Dialog Persona Adherence Evaluates the model’s ability to consistently maintain a specific character or role throughout the conversation.
-Topic adherence Checks if the conversation remains focused on the established topic or transitions logically.
-Linguistic plausibility Assesses whether the dialogue sounds natural, human-like, and contextually appropriate.
-Repetitiveness Measures the degree to which the model avoids unnecessarily repeating phrases or ideas.
-Toxicity check Ensures the response is free from any offensive, harmful, or inappropriate content.
-Instruction adherence Verifies that the model follows meta-instructions given by the user during the dialogue.
-Summarization Coverage Determines if the summary successfully captures all essential points from the source text.
-Factual accuracy Checks that the summary correctly represents the information and facts from the original document.
-Conciseness Evaluates whether the summary is significantly shorter than the source while retaining critical information.
-Coherence and logical flow Assesses if the summary is well-structured, logically organized, and easy to follow.
-Style and tone Measures how well the summary reflects the style and tone of the original text.
-Translation Semantic correctness Assesses whether the meaning, intent, and nuance of the source text are accurately conveyed.
-Fluency correctness Evaluates the grammatical accuracy and naturalness of the translated text in the target language.
-Domain appropriateness Checks if the terminology is correct and suitable for the specific subject matter (e.g., legal, medical).
-Style and tone Determines if the translation successfully captures the original author’s writing style and emotional tone.
-Completeness Verifies that the entire source text has been translated without any omissions or additions.
-A.2 Results from the quality evaluations of the synthetic data
-Figure 3: LLM evaluations across 5 synthetically generated tasks
-Figure 4: Expert human evaluations across 5 synthetically generated tasks
-Figures  3 and 4 show the distribution of the scores received for the tasks we evaluated across the languages for the tasks.
+Table 6: Participant demographics summary.
+Reproducibility
+We provide a detailed reproducibility statement in Appendix A.
+Appendix A Reproducibility Statement
+To ensure full reproducibility of our pipeline—spanning data generation, model fine-tuning, downstream evaluation, and synthetic data quality assessment — we provide our complete codebase and sample generated data in the supplementary material.
+The submission zip contains organized data and code folders, with the code structured into seven components corresponding to distinct pipeline stages:
+  1. 1.
+    selective_translation – Scalable codebase for generating reasoning data through translation from OrcaAgent-Instruct.
+  2. 2.
+    wiki_bt – Wikipedia-based grounded synthetic data curation pipeline.
+  3. 3.
+    DCAD-2000 – Modified repository for heuristic-based filtering of standard and multi-turn conversational data.
+  4. 4.
+    quality_evals – Synthetic data quality evaluation code, including outputs, human annotations, and analysis.
+  5. 5.
+    model_trainingAxolotl-based fine-tuning scripts for the generated datasets.
+  6. 6.
+    lm_evaluation_harness – Custom fork integrating additional benchmarks for downstream evaluation.
+  7. 7.
+    plotting – Analysis and visualization code for experimental results.
+Comprehensive instructions are provided in the main README and component-specific README files for straightforward reproduction.
+Model training configurations with exact hyperparameters are detailed in Appendix 9.
+We will publicly release all the following artifacts (code and datasets) under a permissive license.
+  • Complete codebase including selective translation pipelines, Wikipedia-grounded synthetic data generation, data filtering methods, model training scripts, and the up-to-date evaluation frameworks
+  • Training datasets comprising Updesh Reasoning and Generative subsets across multiple languages.
+  • Evaluation datasets featuring GPT-4o translated variants of standard benchmarks (IFEval, IFBench) for reproducibility and future benchmarking.
+  • Raw evaluation scores for all models across every dataset, providing complete experimental transparency
+  • Human-annotations of synthetic data specifically useful for calibrating / meta-evaluation of LLM evaluators on Indian languages
+Figure 6: Framework for Multilingual & Multicultural Synthetic Data Generation
+Appendix B Appendix
+B.1 Design Considerations
+Our framework (Figure˜6) addresses synthetic data generation for multilingual and multicultural contexts throughout the AI lifecycle — pre-training, SFT, RLHF, and evaluation (Viswanathan et al., 2025).
+Below, we describe the key factors to consider when generating this type of data.
+While our focus is on SFT data, these design considerations can generalize to other synthetic data types.
+Base model capability & Seed data selection
+Select foundation models based on performance in target languages on multilingual benchmarks.
+When language-specific benchmarks are unavailable, use related languages or overall multilingual performance as proxies.
+Other critical aspects to consider include licensing, cost, and model availability (open-weights vs.
+restricted).
+For seed data selection, cover diverse sources and tasks, prioritizing tasks containing cultural knowledge, norms, and values relevant to specific regional contexts.
+Data generation strategy
+Three primary approaches: (i) Translation from English SFT datasets to transfer critical skills, though risking translationese artifacts (Zhang and Toral, 2019; Vanmassenhove et al., 2021); (ii) Back-translation—using existing unlabeled multilingual datasets through instruction backtranslation (Li et al., 2024) or back-and-forth translation (Nguyen et al., 2024); (iii) Retrieval-augmented generation —leveraging curated native speaker-authored content from web corpora to capture cultural knowledge and linguistic nuances.
+Translation-based approaches yield weaker correlations with human judgements than language-specific benchmarks (Kreutzer et al., 2025; Wu et al., 2025).
+Bottom-up approaches grounded in web corpora show superior performance (Shaham et al., 2024; Khan et al., 2024; Doshi et al., 2024) but remain underexplored in multilingual contexts.
+Quality metrics
+Essential dimensions for assessment: Language correctness—proper language, register, dialect identification (Marchisio et al., 2024); Linguistic acceptability—native speaker fluency and naturalness (Hada et al., 2024b, a); Cultural appropriateness—accurate cultural references, values, and norms; Bias and safety—absence of stereotypes and culturally inappropriate content (Pawar et al., 2025).
+Downstream evaluation
+Select benchmarks covering all target languages, avoiding English-translated datasets.
+Include diverse tasks testing cultural knowledge and values.
+Address benchmark contamination risks (Ahuja et al., 2024) and develop new benchmarks when necessary.
+Native speaker involvement
+Engage native speakers in seed selection and evaluation.
+Ensure informed consent addressing cultural considerations and data sovereignty per local regulations.
+Exclude personally identifiable information from all sources.
+B.2 Task Descriptions
+Task Descriptions for all the subtasks in Updesh can be found in Table 7 and 8.
+Task Type Description
+Analytical Reasoning MCQ-style questions requiring step-by-step logical inference
+Multiple-Choice Questions General-purpose problems across diverse knowledge domains
+Fermi (Guesstimation) Open-ended estimation problems using logical assumptions
+Few-Shot Chain-of-Thought Tasks with 4-5 in-context examples for learning
+Brain Teasers Puzzles stimulating lateral thinking and creativity
+Text Classification Categorization tasks for predefined labels
+Reading Comprehension Questions based on understanding and interpreting textual passages
+Math Grade-school arithmetic, algebra, and geometry word problems
+Table 7: Reasoning task categories and their descriptions.
+Task Type Synthesis Method Phases Qwen3-Mode
+Logical Reasoning Generate implicit inferences from text passages (1) Direct inference generation Reasoning
+Multi-Hop QA Create questions requiring information synthesis across text segments (1) Question generation (2) Answer generation Reasoning
+Creative Writing Transform factual content into engaging narratives (1) Generate creative piece (2) Generate eliciting prompt Reasoning
+Multi-Turn Dialogue Agentic workflows with 3-5 turn conversations between personas (1) Generate dialog adhering to personas (2) Generate natural prompt Non-reasoning
+Summarization Generate summaries preserving key information across languages (1) Direct summary generation Non-reasoning
+Machine Translation Cross-lingual conversion maintaining cultural context (1) Direct translation Non-reasoning
+Causal Reasoning Identify and explain cause-effect relationships in text (1) Direct causal analysis Reasoning
+Table 8: Generative task categories with synthesis methods, phases, and model configuration
+B.3 Rubrics used for the quality evaluation of the synthetic data
+Creative Writing
+  • Instruction adherence: Assesses if the output strictly follows all constraints and guidelines provided in the prompt.
+  • Fluency: Evaluates the naturalness, grammatical correctness, and readability of the generated text.
+  • Narrative coherence: Checks for logical consistency in the plot, character development, and thematic elements.
+Reasoning Tasks
+  • Answer adequacy: Determines if the final answer is correct, complete, and directly addresses the core question.
+  • Context adherence: Measures whether the reasoning remains faithful to the provided context, avoiding external facts.
+  • Instruction adherence: Verifies that the output’s structure, format, and steps match the user’s instructions.
+  • Fluency and readability: Assesses the clarity, logical flow, and ease of understanding of the explanation.
+  • Problematic content and cultural relevance: Scrutinizes the response for harmful stereotypes and ensures it is culturally sensitive and appropriate.
+Multi-turn Dialog
+  • Persona adherence: Evaluates the model’s ability to consistently maintain a specific character or role throughout the conversation.
+  • Topic adherence: Checks if the conversation remains focused on the established topic or transitions logically.
+  • Linguistic plausibility: Assesses whether the dialogue sounds natural, human-like, and contextually appropriate.
+  • Repetitiveness: Measures the degree to which the model avoids unnecessarily repeating phrases or ideas.
+  • Toxicity check: Ensures the response is free from any offensive, harmful, or inappropriate content.
+  • Instruction adherence: Verifies that the model follows meta-instructions given by the user during the dialogue.
+Summarization
+  • Coverage: Determines if the summary successfully captures all essential points from the source text.
+  • Factual accuracy: Checks that the summary correctly represents the information and facts from the original document.
+  • Conciseness: Evaluates whether the summary is significantly shorter than the source while retaining critical information.
+  • Coherence and logical flow: Assesses if the summary is well-structured, logically organized, and easy to follow.
+  • Style and tone: Measures how well the summary reflects the style and tone of the original text.
+Translation
+  • Semantic correctness: Assesses whether the meaning, intent, and nuance of the source text are accurately conveyed.
+  • Fluency correctness: Evaluates the grammatical accuracy and naturalness of the translated text in the target language.
+  • Domain appropriateness: Checks if the terminology is correct and suitable for the specific subject matter (e.g., legal, medical).
+  • Style and tone: Determines if the translation successfully captures the original author’s writing style and emotional tone.
+  • Completeness: Verifies that the entire source text has been translated without any omissions or additions.
+B.4 Results from the quality evaluations of the synthetic data
+Figures  7 and 8 show the distribution of the scores received for the tasks we evaluated across the languages for the tasks.
 Expert human evaluators have consistently given a score of 2 across languages and tasks, indicating the high quality of Updesh.
@@
 We thereby, performed a thorough inter annotator analysis, the details of which are present in the next figures.
-Figure 5: Confusion matrices showing agreement between human and LLM evaluators
-Figure 6: Agreement between human and LLM evaluators per task and language respectively
 Continuing the claims made from Figure 2, Figure 9 provides a clearer view of the tasks on which humans and LLMs are likely to agree or disagree.
@@
 LLMs also perform reasonably well at identifying whether a text is culturally relevant, but issues arise when the evaluation requires more fine-grained judgments of multilinguality and multiculturalism.
-We do not identify any specific trends language or task-wise as apparent in Figure6.
-A.3 Data Quality Assesment of Reasoning Data
-Language analytical brain fermi fs (cot_flow) math mcq rc text_class
+We do not identify any specific trends language or task-wise as apparent in Figure˜10.
+Figure 7: LLM evaluations across 5 synthetically generated tasks
+Figure 8: Expert human evaluations across 5 synthetically generated tasks
+Figure 9: Confusion matrices showing agreement between human and LLM evaluators
+Figure 10: Agreement between human and LLM evaluators per task and language, respectively
+B.5 Generation Hyperparameters
+For data synthesis, decoding is performed using nucleus sampling with top_p=0.95top\_p=0.95 and $temperature=1.0$.
+B.6 Training Hyperparameters
+Hyperparameters for all our training runs could be found in Table 9
+Hyperparameter Value
+Base Model phi4-base / llama-3.1-8b
+Sequence Length 65,536
+Effective Batch Size 8192
+Number of Epochs 3
+Optimizer AdamW
+Learning Rate 1e-5
+LR Scheduler Cosine
+Adam Betas (0.9, 0.95)
+Max Grad Norm 1.0
+Warmup Ratio 0.03
+Weight Decay 0.1
+NEFTune Noise Alpha Jain et al. (2024) 5
+Precision BF16
+Flash Attention True
+Gradient Checkpointing True
+Table 9: Training hyperparameters used for all our experiments.
+B.7 Data Quality Assesment of Reasoning Data
+Language Analytical Brain Fermi Fs-CoT Math MCQ RC Text Class
 Assamese 75.02 71.62 79.93 96.59 79.87 64.33 65.07 71.29
@@
 Table 10: Backtranslation ChrF scores for the Reasoning subset.
-A.4 Hyperparameters
-Table 10: Training hyperparameters used for all our experiments
-Hyperparameter Value
-Base Model phi4-base-hf / llama-3.1-8b-hf
-Sequence Length 65,536
-Effective Batch Size 8192
-Number of Epochs 3
-Optimizer AdamW Torch Fused
-Learning Rate 1.0e-05
-LR Scheduler Cosine
-Adam Beta1 0.9
-Adam Beta2 0.95
-Max Grad Norm 1.0
-Warmup Ratio 0.03
-Weight Decay 0.1
-NEFTune Noise Alpha 5
-Precision BF16
-Flash Attention True
-Gradient Checkpointing True
-A.5 Evaluation Datasets
-Table 11: Evaluation datasets
-Dataset Source
-NLU Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021)
-MMLU Indic (MMLU-I) SarvamAI
-ARC Indic (ARC-I) SarvamAI
-BoolQ Indic (BoolQ-I) SarvamAI
-TriviaQA Indic (TVQA-I) SarvamAI
-BeleBele (Bele) Bandarkar et al. (2024)
-INCLUDE (INCL) Romanou et al. (2025)
-Global MMLU (GMMLU) Singh et al. (2025)
-NLG Extreme Summarization (Xsum) Singh et al. (2024a)
-Flores English to Other (Flores EnXX) Goyal et al. (2022)
-Flores XX-En (Flores XXEn)
-IN22-Conv (in22-conv-doc) Gala et al. (2023); Gumma et al. (2025)
-IF Multilingual Instruction Following Evaluation (mIFEval) Zhou et al. (2023b)
-Multilingual Instruction Following Benchmark (mIFBench) Pyatkin et al. (2025)
-A.6 Prompt used for conversion of IFEval and IFBench
-Translation Generation Prompt INSTRUCTION: You are a professional translator.
-Your sole task is to translate the provided English text into {target_language}.
-Rules: Output only the translation—nothing else (no explanations, no commentary).
-Do not answer or solve any task contained in the text; just translate it.
-Preserve formatting, punctuation, numbers, placeholders, and bracketed tokens exactly as they appear.
-Translate idiomatically and naturally for a fluent native speaker of {target_language}.
-If the input is an instruction (e.g., “Write a resume for …”), translate the instruction itself—do not perform it.
-Do not introduce extra formatting (e.g., don’t create JSON blocks unless the source already has them).
-SOURCE LANGUAGE: English TARGET LANGUAGE:
-{target_language}
-TEXT TO TRANSLATE:
-{source_text}
-OUTPUT REQUIREMENT:
-Return only the translated text in {target_language} with formatting faithfully preserved.
-A.7 Sample Prompts Used for Evaluation
-The following are representative prompts (one per task) used to evaluate different task types. Each prompt follows the established rubrics from Table LABEL:tab:rubrics.
-A.7.1 Creative Writing - Instruction Adherence
+B.8 Detailed Results
+Model Setting NLU Avg MMLU-I MILU ARC-I BoolQ-I TVQA-I Bele INCL GMMLU
+Llama 8B Base 29.33 22.97 25.88 25.26 62.04 26.55 24.06 24.88 23.02
+Bactrian-X 45.37 36.80 39.41 26.49 71.21 66.30 41.45 39.14 42.14
+IndicAlign 42.14 34.47 39.67 27.35 71.14 59.35 32.09 34.37 38.67
+Aya-Collection 42.73 32.69 36.98 28.11 76.24 50.65 46.54 33.84 36.81
+Updesh 42.89 34.40 36.13 29.67 77.51 46.13 51.30 30.56 37.43
+Phi-4 Base 53.17 43.38 49.22 29.99 79.56 68.25 54.55 45.46 54.97
+Bactrian-X 49.25 39.91 45.09 29.40 74.40 67.30 48.32 41.72 47.89
+IndicAlign 51.71 44.99 48.25 32.23 76.26 61.68 51.68 44.71 53.92
+Aya-Collection 54.41 45.08 50.57 33.01 72.62 71.80 60.71 48.02 53.47
+Updesh 56.61 48.98 51.19 32.08 81.38 60.14 74.42 47.24 57.47
+Table 11: Performance comparison of Llama 8B and Phi-4 variants across Indic NLU Tasks. All entries are accuracy (higher is better)
+Model Setting NLG Avg Flores En-XX Flores XX-En XSum IN22-Conv-Doc En-XX IN22-Conv-Doc XX-En
+Llama 8B Base 8.91 1.45 41.71 0.16 0.60 0.60
+Bactrian-X 23.81 28.85 50.98 0.21 19.51 19.51
+IndicAlign 15.67 32.12 3.20 12.84 15.09 15.09
+Aya-Collection 9.69 28.38 0.46 0.23 9.69 9.69
+Updesh 36.41 44.00 51.88 25.54 30.31 30.31
+Phi4 Base 34.20 30.23 56.57 17.59 33.31 33.31
+Bactrian-X 24.45 26.78 51.86 0.31 21.60 21.71
+IndicAlign 13.77 32.13 0.59 0.28 17.89 17.94
+Aya-Collection 13.93 30.16 1.58 0.37 18.77 18.78
+Updesh 36.56 45.82 59.55 21.66 27.81 27.94
+Table 12: Performance comparison of Llama 8B and Phi-4 variants across Indic NLG. All entries are ChrF scores (higher is better).
+Model Variant Avg as bn en gu hi kn ml mr ne or pa ta te ur
+Llama 8B Base ZS 28.98 22.22 29.61 36.08 29.97 29.19 30.41 30.82 30.72 23.02 31.14 30.14 29.91 28.97 23.44
+Bactrian 45.65 38.22 45.73 68.19 44.06 50.88 43.41 44.07 48.05 40.12 40.24 44.90 40.81 40.35 50.11
+IndicAlign 41.03 32.11 42.04 63.79 42.05 44.55 42.36 41.30 43.38 31.84 38.87 41.52 40.79 39.83 30.00
+Aya 44.30 42.78 42.43 63.04 42.07 45.66 43.28 43.76 43.15 36.03 38.52 43.06 41.74 40.85 53.89
+Updesh 45.23 46.44 41.97 59.84 44.52 42.35 44.95 45.56 42.99 35.63 43.99 44.08 41.82 41.98 57.11
+Phi-4 Base ZS 53.36 42.22 52.77 75.64 53.93 60.37 51.55 49.40 53.25 48.14 47.51 54.21 48.00 47.84 62.22
+Bactrian 50.16 40.67 49.39 77.47 49.21 56.36 45.57 45.53 50.54 45.06 44.47 50.60 41.87 41.80 63.67
+IndicAlign 52.86 49.33 51.13 75.84 50.74 55.49 50.33 50.88 49.38 49.91 42.72 51.22 48.07 47.44 67.56
+Aya 55.80 52.89 53.41 76.41 55.16 60.23 54.50 51.63 55.01 52.87 46.25 55.46 50.10 49.33 68.00
+Updesh 59.71 70.33 55.77 75.45 56.91 61.82 56.43 54.20 56.81 58.21 53.35 57.59 51.49 52.04 75.56
+Table 13: NLU Performance: Evaluation across Indic languages including English.
+Language codes: as: Assamese, bn: Bengali, en: English, gu: Gujarati, hi: Hindi, kn: Kannada, ml: Malayalam, mr: Marathi, ne: Nepali, or: Odia, pa: Punjabi, ta: Tamil, te: Telugu, ur: Urdu.
+Model Variant Avg as bn gu hi kn ml mr ne or pa ta te ur
+Llama 8B Base ZS 9.08 7.05 7.87 7.76 11.05 9.22 9.04 9.77 11.83 8.48 10.26 7.30 9.24 9.22
+Bactrian 24.24 20.32 29.79 24.24 33.13 22.87 23.97 19.58 28.30 18.26 21.63 24.24 23.32 25.51
+IndicAlign 15.80 13.09 18.29 12.27 26.29 15.34 14.75 17.43 21.57 6.65 13.71 17.22 14.66 14.14
+Aya 9.87 7.49 12.98 10.83 8.62 11.07 12.74 8.04 11.86 6.42 8.34 12.00 8.78 9.10
+Updesh 36.68 33.80 35.98 34.90 36.81 35.73 37.75 34.15 43.13 32.87 35.60 39.08 39.64 37.42
+Phi-4 Base ZS 34.56 25.39 39.47 38.83 46.31 30.02 30.97 32.36 40.76 27.36 40.32 30.73 34.61 32.10
+Bactrian 24.96 20.74 30.75 24.90 40.36 22.72 23.68 23.53 33.54 13.29 17.42 25.19 25.27 23.14
+IndicAlign 14.08 12.81 16.57 15.35 25.27 12.46 11.59 13.06 20.42 6.95 10.81 15.58 19.52 2.60
+Aya 14.29 7.67 21.88 15.44 25.36 5.15 12.57 12.80 23.35 1.91 16.39 12.33 15.49 15.40
+Updesh 36.87 32.60 37.66 38.50 40.38 34.21 36.50 35.12 42.16 33.56 36.29 39.99 37.61 34.74
+Table 14: NLG Performance: Evaluation across Indic languages (excluding English).
+Language codes: as: Assamese, bn: Bengali, gu: Gujarati, hi: Hindi, kn: Kannada, ml: Malayalam, mr: Marathi, ne: Nepali, or: Odia, pa: Punjabi, ta: Tamil, te: Telugu, ur: Urdu.
+Model Variant Avg awa bho brx hne gbm bgc gom mai mup mni mwr ps hoj sa sat bo
+Llama 8B Base ZS 10.13 15.86 14.82 2.64 13.06 13.72 15.41 7.24 9.34 13.66 1.00 15.13 13.29 13.88 5.81 2.80 4.39
+Bactrian 19.59 27.70 19.28 4.45 28.41 27.56 26.20 13.23 23.72 26.71 5.00 29.27 20.09 26.89 17.38 4.25 13.29
+IndicAlign 9.19 12.45 11.03 0.32 15.52 11.10 15.35 6.00 10.16 13.79 2.23 14.36 7.14 14.43 11.59 0.21 1.38
+Aya 6.10 6.97 7.63 0.24 7.28 8.87 9.23 4.28 7.16 8.93 2.16 9.36 3.16 9.78 6.55 0.23 5.82
+Updesh 24.40 33.67 32.33 4.30 37.34 30.20 35.61 23.53 26.23 33.13 6.60 36.50 21.46 33.69 21.25 4.55 10.04
+Phi-4 Base ZS 22.69 35.45 32.75 5.01 30.05 31.04 26.78 15.74 31.99 27.04 5.73 29.62 25.09 24.09 26.08 4.99 11.61
+Bactrian 22.19 31.71 27.96 4.80 31.14 30.38 31.70 15.38 26.43 30.58 5.90 31.13 19.33 30.11 21.13 3.80 13.64
+IndicAlign 9.41 12.51 10.26 1.33 12.83 12.65 12.83 9.09 17.01 13.51 1.28 13.52 6.76 10.14 11.03 2.21 3.58
+Aya 12.20 17.67 13.22 1.12 16.44 16.89 17.25 7.98 18.41 14.90 0.55 16.13 12.47 14.14 14.86 4.25 8.95
+Updesh 23.51 34.32 30.71 4.90 35.04 32.39 28.16 21.88 26.82 31.22 7.29 33.82 21.04 29.12 22.06 4.19 13.15
+Table 15: Detailed performance breakdown for out-of-training languages.
+Language codes: awa: Awadhi, bho: Bhojpuri, brx: Bodo, hne: Chhattisgarhi, gbm: Garhwali, bgc: Haryanvi, gom: Konkani, mai: Maithili, mup: Malvi, mni: Manipuri, mwr: Marwari, ps: Pashto, hoj: Rajasthani, sa: Sanskrit, sat: Santali, bo: Tibetan.
+Figure 11: Scatter plots correlating ELO Scores with NLU Average (left), NLG Average (center), and Dataset Wins (right).
+The Updesh-32K model (UB 32K) consistently outperforms baselines, appearing in the top-right quadrant across all metrics.
+Appendix C Cultural Evaluation Framework
+To assess the local cultural and community-specific knowledge of Large Language Models (LLMs) within the Indian context, we collected a set of questions in collaboration with a third-party non-profit organization.
+The dataset consists of 4,399 unique queries that reflect authentic information-seeking behaviors of local populations.
+These queries span 11 Indian languages (see Figure 12) and cover four high-impact domains: Healthcare, Education, Finance, and Legal.
+To ensure a robust assessment, we performed a comparative evaluation of all model checkpoints, totaling 91,982 pairwise comparisons (battles).
+We refer the reader to Hamna et al.
+(2025) for the collection process, which we replicated.
+Domain Themes.
+The dataset covers a wide spectrum of community needs and themes, some of which were:
+  • Education: Teaching and learning support, and career guidance.
+  • Finance: Insurance, savings, and budgeting strategies.
+  • Healthcare: Senior care protocols and general wellness habits.
+  • Legal: Product/service disputes and family or marriage-related inquiries.
+Figure 12: Cultural evaluation language wise distribution
+Figure 13: Exanples of a Punjabi and Malayalam query we collected
+C.1 Sample Prompts Used for Evaluation
+The following are representative prompts (one per task) used to evaluate different task types. Each prompt follows the established rubrics from Section B.3.
+C.1.1 Creative Writing - Instruction Adherence
 Creative Writing Instruction Adherence Evaluation Prompt INSTRUCTION: You are an expert literary critic and evaluator, tasked with assessing the degree to which a synthetically generated creative piece adheres to the user’s writing prompt.
@@
   "score": <integer from 1 to 5>,
-  "reason": "<brief explanation for the score>"
+  "reason": "<brief explanation>"
 Do not include markdown, comments, or anything outside the JSON.
-A.7.2 Multi-turn Dialog - Persona Adherence
+C.1.2 Multi-turn Dialog - Persona Adherence
 Multi-turn Dialog Evaluation Prompt INSTRUCTION: You are an expert evaluator tasked with assessing how well a multi-turn dialog maintains persona consistency.
@@
   "score": <integer from 1 to 5>,
-  "reason": "<brief explanation for the score>"
+  "reason": "<brief explanation>"
 Do not include markdown, comments, or anything outside the JSON.
-A.7.3 Reasoning Tasks - Answer Adequacy
+C.1.3 Reasoning Tasks - Answer Adequacy
 Reasoning Task Evaluation Prompt INSTRUCTION: You are an expert evaluator specializing in logical reasoning and problem-solving.
@@
   "score": <integer from 1 to 5>,
-  "reason": "<brief explanation for the score>"
+  "reason": "<brief explanation>"
 Do not include markdown, comments, or anything outside the JSON.